### v1

# Cyber Defense Reinvented: Large Language Models as Threat Intelligence Copilots

Abstract
The exponential growth of cyber threat knowledge, exemplified by the expansion of databases such as MITRE-CVE and NVD, poses significant challenges for cyber threat analysis. Security professionals are increasingly burdened by the sheer volume and complexity of information, creating an urgent need for effective tools to navigate, synthesize, and act on large-scale data to counter evolving threats proactively. However, conventional threat intelligence tools often fail to scale with the dynamic nature of this data and lack the adaptability to support diverse threat intelligence tasks.
In this work, we introduce CyLens, a cyber threat intelligence copilot powered by large language models (LLMs). CyLens is designed to assist security professionals throughout the entire threat management lifecycle, supporting threat attribution, contextualization, detection, correlation, prioritization, and remediation. To ensure domain expertise, CyLens integrates knowledge from 271,570 threat reports into its model parameters and incorporates six specialized NLP modules to enhance reasoning capabilities. Furthermore, CyLens can be customized to meet the unique needs of different organizations, underscoring its adaptability. Through extensive evaluations, we demonstrate that CyLens consistently outperforms industry-leading LLMs and state-of-the-art cybersecurity agents. By detailing its design, development, and evaluation, this work provides a blueprint for leveraging LLMs to address complex, data-intensive cybersecurity challenges.
1 Introduction
We are living in an era of rapid digital transformation, where technological advancements are closely linked to the growing prevalence of cyber threats. In recent years, the landscape of cyber threats has evolved dramatically, marked by a 25% annual increase in reported CVEs [40]. The increase in CVEs can be attributed to several factors, especially the growing complexity of IT systems [76], the widespread adoption of open-source software [22, 79], and the accelerated pace of development cycles [83]. Together, these factors facilitate the exploitation of cyber threats in modern software ecosystems.
To enable efficiently analyze and counteract these threats, Cyber Threat Intelligence (CTI) [23] has emerged as a crucial strategy. CTI provides actionable insights derived from the analysis of threat data and empowers security teams to anticipate, identify, and mitigate cyber risks effectively [89, 110, 8].
However, CTI faces critical challenges in scalability and adaptability. First, the overwhelming volume of threat data creates significant bottlenecks [1]. Security teams are often inefficiently processing and correlating this massive influx of security knowledge, limiting their ability to respond to emerging cyber threats. Second, integrating CTI into existing cybersecurity platforms often requires extensive customization, which poses a significant barrier for small and medium-sized organizations who lack resources to deploy CTI effectively. To address these challenges, it is essential to propose advanced CTI solutions capable of scalably collecting and analyzing extensive cybersecurity knowledge while adaptively support customization tailored to diverse organizational needs.
Figure 1: (a) The lifecycle of threat management. Each task consists of multiple analytical targets, as detailed in Table 1. (b) The development methodology of CyLens, as detailed in section 3 and section 4.
Large language models (LLMs) have recently demonstrated exceptional capabilities, driven by their scalability to massive knowledge through pre-training and their adaptability for organization-specific applications via fine-tuning. Beyond these strengths, LLMs can be further augmented with task-oriented prompts or external modules, enabling more advanced analyses that require topic modeling [96], entity and relation recognition [97, 37], and retrieval-augmented generation (RAG) [33, 10]. Hence, LLMs emerge as a promising solution for assisting humans in specialized fields such as bug fixing [72, 94] and fuzz testing [38, 104, 24]. Consequently, we raise a critical question: how can LLMs be systematically developed as copilots to support CTI activities?
This work. To address this question, we propose CyLens, an LLM-powered copilot system designed to enhance CTI. CyLens consists of several sizes—1B, 8B, and 70B—to accommodate diverse needs and computational constraints. We characterize CyLens by the following aspects:
Lifecycle-Oriented Threat Management – As illustrated in Figure 1-(a), CyLens supports six tasks that cover the entire lifecycle of threat management. To analyze the properties of threats, CyLens enables ❶ attribution by identifying threat actors and gathering technical evidence, such as TTPs and exploitation paths, to trace the origin of a threat. It also facilitates ❷ contextualization by enriching metadata—such as affected systems—to provide a more comprehensive understanding of the threat landscape. Next, ❸ detection involves specifying available identifiers, such as CVE [14], while ❹ correlation enables CyLens to uncover adversarial networks by identifying relationships between the current threat and others. CyLens then drives ❺ prioritization by quantifying the severity and predicting the likelihood of exploitation. Building on this analytical foundation, ❻ remediation effectively counteracts threats and limits their broader impact.
Collective Expertise. As shown in Figure 1-(b), CyLens manages web-scale CTI knowledge through a combination of development methodologies. Our approach is motivated by the fact that cyber threats co-exist as an interdependent community, where attack vectors, infrastructure, and exploitations are deeply interconnected [46, 6, 93, 12, 103]. Effectively addressing these complex relationships requires a vast, integrated set of expertise rather than isolated knowledge.
To achieve this, CyLens first undergoes pre-training on a diverse set of databases—MITRE-CVE [14], CWE [16], CAPEC [17], ATT&CK [18], NVD [59], Exploit-DB [62], and correlated third-party repositories (e.g., [39, 5])—to incorporate extensive CTI knowledge from 271,570 threat reports. We introduce a curriculum pre-training approach, where “curriculum” data is designed to progressively embed CTI expertise into CyLens. Following pre-training, CyLens undergoes fine-tuning using task-oriented datasets addressing critical CTI tasks (❶-❻). To enhance its analytical depth, we employ a cascading reasoning process, allowing CyLens to systematically build awareness of dependencies within CTI analysis. Additionally, we integrate specialized design modules to reinforce CyLens ’s ability to conduct rigorous reasoning, escalating it as a human-usable CTI copilot.
Operational Proficiency. Our experiments demonstrate CyLens’s effectiveness across all above tasks under two categories of cyber threats: (i) Historical Threats—previously recorded threats that security analysts may not be fully aware of, and (ii) Zero-Day Threats—newly emerging threats that have not yet been documented. Across both threat categories, CyLens consistently outperforms SOTA models and industry-leading LLMs (e.g., ChatGPT and Gemini), underscoring its versatility in handling diverse cybersecurity challenges while excelling in proactive threat hunting within the ever-evolving cyberattack landscape.
Adaptive to Customization. CyLens also demonstrates adaptability in addressing organization-specific threats related to particular products, vendors, or threat groups. As shown in our evaluations in section 5.6, CyLens consistently delivers superior and balanced performance across diverse organizational requirements. This adaptability is particularly valuable when an organization’s threat-related data volume is insufficient, enabling effective CTI analysis even in data-scarce scenarios.
Contributions. We summarize following contributions:
    • LLM-as-CTI-Copilot: We propose CyLens, a scalable and adaptive system for CTI practice. CyLens is designed to accommodate diverse needs, making it well-suited for both research and industry applications.
    • Empirical Insights: Through extensive evaluations, we not only demonstrate CyLens’s effectiveness, but also reveal the varying performance of SOTA and industry-leading LLMs, highlighting the inherent complexity of CTI analysis and difficulty of achieving consistent expertise.
    • Blueprint Perspective: The design and development of CyLens serve as a blueprint for integrating LLM copilots into cybersecurity practice. This endeavor not only advances the capabilities of CTI systems but also provides insights that could drive the development of expert-level AI across other high-stakes domains [^1] Codes and datasets are organized at https://anonymous.4open.science/r/security-agent-4BB7/.
2 Background and Related Work
2.1 Cyber Threat Ingelligence (CTI)
AI-for-Security has been a long-standing research area, with extensive studies covering a broad range of cybersecurity topics, including anomaly detection [109, 68], malware classification[47, 69], and intrusion detection[48, 31]. Among these, CTI has emerged as a systematic approach to threat management, enabling organizations to collect, analyze, and operationalize intelligence on cyber threats [20, 95, 89].
Although there is no universal specification for CTI practices, we categorize CTI into six critical threat management tasks based on extensive literature [82, 32, 86, 54, 42, 75, 89, 71, 114, 43] and the MITRE initiative [3], a foundational effort in the field. As illustrated in Figure 1-(a), the lifecycle of threat intelligence begins with the extraction and analysis of threat properties (❶ & ❷) [42, 105, 75], progresses to the identification of adversary community (❸ & ❹) [32, 86, 54, 114], and culminates in the development of effective countermeasures (❺ & ❻) [89, 71, 43]. In this work, we set the scope of the CTI tasks as follows:
    • Attribution–Identifying the origin of cyber threats by profiling threat actors, analyzing tactics, techniques, and procedures (TTPs), and tracing exploitation paths to link cyber incidents to specific adversaries or groups. Attribution helps organizations understand attacker intent, capabilities, and historical attack patterns [105, 42].
    • Contextualization–Enriching raw threat intelligence with relevant metadata, such as affected systems, attack infrastructures, and potential impacts. Contextualization enables security analysts to assess threats with a comprehensive picture of the threat landscape[52, 75].
    • Detection–Identifying cyber threats using available indicators, such as CVE [18] and CWE [16]. Detection assists the recognition of both known and emerging threats [114, 32].
    • Correlation–Connecting disparate threats to uncover hidden relationships between attack campaigns and disclose the broader landscape of adversarial community. Correlation enhances situational awareness and strengthens incident investigations by providing a more comprehensive view of threat activities [86, 54, 52].
    • Prioritization– Ranking threats based on severity, exploitability, and potential impact to determine the urgency of remediation. Effective prioritization optimizes resource allocation by focusing on the most critical threats [89, 43].
    • Remediation–Developing and recommending actionable countermeasures, including security patches, mitigation strategies, and responsible advisories. Remediation ensures that organizations can effectively neutralize threats and reinforce their security posture[43, 74].
Within each CTI task, we define distinct analytical targets. For instance, prioritization involves assessing both severity, measured by the Common Vulnerability Scoring System (CVSS) [60], and exploitation potential, measured by the Exploit Prediction Scoring System (EPSS) [29]. The detailed analytical targets for each CTI task are presented in Table 1 and are evaluated extensively in section 5.
2.2 LLM-as-Copilot
Figure 2: Development pipeline of LLM copilots.
Due to their vast inherent knowledge and ability to handle large-scale downstream tasks, LLMs have been extensively adapted for domain-specific applications, an effort referred to as “LLM-as-Copilot” [51, 34, 90]. Unlike the broader concept of an "agent," which is designed to perform human-like behaviors, a copilot specifically emphasizes its role as a human assistant [85, 55, 88]. This distinction is particularly critical in high-stakes domains such as cybersecurity [112, 28], healthcare [4, 98, 108], and finance [107, 111, 70], where expert guidance and decision support are essential.
Training Methodology. As illustrated in Figure 2, the development of LLM copilots follows a pre-training and fine-tuning trajectory. Pre-training leverages large-scale corpora to embed extensive domain expertise into base LLMs through a causal language modeling objective [9, 25, 45, 27]. Subsequently, fine-tuning incorporates environment-aware instructions and utilizes supervised learning [61, 78] to refine the model’s reasoning capabilities for practical applications.
This pre-training followed by fine-tuning approach has proven particularly effective in agent development requiring broad domain expertise, such as existing efforts in AI-assisted programming [2, 91]. Similarly, the interconnected nature of cyber threats necessitates a deep and expansive CTI knowledge. Thus, this work adopts the same development pipeline to enhance LLM capabilities, advancing them toward the creation of an effective and practical cybersecurity copilot.
Prompting. Besides training methodologies, inference-time prompting plays a crucial role in regulating an LLM’s behavior during execution [63, 36]. While developing LLM copilots cannot rely solely on prompting, it provides flexibility by enabling them to interact with real-world resources, such as functional tools [87] or retrieving web-based knowledge [53].
In this work, we enhance CyLens’s inference by implementing 6 NLP modules, integrated through either additional functions or strategic prompting techniques. These enhancements improve CyLens’s reasoning, contextual awareness, and overall effectiveness in CTI analysis.
3 Data Management and Pre-Training
In this section, we describe the collection and preprocessing of threat reports used in pre-training (section 3.1) and introduce a curriculum-based pre-training approach that integrates large-scale cyber threat knowledge into CyLens (section 3.2). The subsequent section (section 4) details the fine-tuning and inference design.
3.1 Threat Data: Collection and Preprocessing
Raw Data Collection. Given that threat data is predominantly represented by CVEs [100, 102], we focus on collecting CVE-centered metadata from several vulnerability databases: MITRE-CVE [14], NVD [59], Exploit-DB [62], CWE [16], CAPEC [17], and ATT&CK [18], including (but not limited to) associated threat actors, TTPs, affected systems, attack infrastructures, and strategical mitigation approaches.
To enrich the collected cyber threat knowledge, we further extract reports and evidence from third-party references such as GitHub commits, Apache [5], and IBM X-Force [39], incorporating diverse cybersecurity activities and countermeasure strategies. These additional sources enhance the comprehensiveness of our dataset by providing real-world exploit details and remediation insights. Table 8 outlines the data sources and the corresponding metadata collected.
Pre-Training Corpus. Next, we preprocess the collected metadata to construct CVE-centric threat reports, which are used as pre-training corpus. As outlined in Algorithm 1, we employ a combination of three approaches to ensure the generation of high-quality reports and diverse layouts:
    • In-Context Learning: We extract sample layouts and phrasing styles from threat reports published by Microsoft [13], Cisco [41], CISA [20], Europol [26], and ENISA [30]. These examples are used as demonstrations $\mathcal{S}$ to guide the generation. For each generation, we select an example report $s\in\mathcal{S}$ and construct a 1-shot prompt, designed to replicate the specific style and layout of $s$.
    • Multi-Agent Generation: For each report generation, we randomly select one LLM from a model set $\Theta$, including GPT-4o, GPT-o1, Gemini-Pro, Pixtral-Large, DeepSeek-R1, or Llama-3.1-405B, to generate threat reports using the collected metadata and a 1-shot prompt. Subsequently, a different LLM is randomly chosen from $\Theta$ to revise the generated reports, introducing variability and enhancing quality by reducing single-model bias.
    • Output Sampling: We adjust two decoding hyperparameters—temperature and top-p—which introduce controlled variability in the outputs. This approach, as studied in [11, 72, 115], helps balance randomness and coherence, ensuring diverse yet meaningful generations.
Input: $\Theta$ – a set of LLMs; $\mathcal{M}$ – collected metadata; $\mathcal{S}$ – example threat reports; $\mathcal{P}_{\texttt{gen}}$ – generation prompt; $\mathcal{P}_{\texttt{rev}}$ – revision prompt;
Output: $D$ – cyber threat corpus;
1 $D\leftarrow\emptyset$ ;
2 foreach metadata for a CVE: $m\in\mathcal{M}$ do
       // Prompt Construction
3       1-shot demo: $s\leftarrow\texttt{RandSelectFrom}(S)$ ;
4       Input prompt: $\mathcal{P}_{m}\leftarrow\mathcal{P}\oplus s\oplus m$;
       // Report Generation and Revision
5       $\theta_{1},\theta_{2}\leftarrow\texttt{RandSelectFrom}(\Theta)$ ;
6       Randomly set temperature $\mathcal{T}_{1},\mathcal{T}_{2}$ and top-p $p_{1},p_{2}$;
7
8      Generation: $d\leftarrow\texttt{OutSampling}(\theta_{1},\mathcal{P}_{m},\mathcal{T}_{1},p_{1})$ ;
9       Input prompt: $\mathcal{P}_{m}\leftarrow\mathcal{P}_{\texttt{rev}}\oplus d$ ;
10       Revision: $d\leftarrow\texttt{OutSampling}(\theta_{2},\mathcal{P}_{m},\mathcal{T}_{2},p_{2})$ ;
11
12      $D\leftarrow D\cup\{d\}$;
13
14 end foreach
15return $D$;
Algorithm 1 Corpus Generation
Due to space constraints, we defer the example reports, input prompts, and the corpus scale to Appendix A.2.
Table 1: Details of the fine-tuning dataset. Each task targets a specific aspect of CTI analysis, incorporating varying dependencies from prior tasks, and involving different NLP modules: TOM–Topic Modeling; NER–Named Entity Recognition; REL–Relation Extraction; RAG–Retrieval-Augmented Generation (using web resources); REA–(any type of) Reasoning; SUM–Summarization.
CTI Task Analytical Target Dependency NLP Module
TOM NER REL RAG REA SUM
❶ Attribution Threat Actor, TTPs, Exploitation Path Initial Evidence
❷ Contextualization Affected System, Attack Infra, Impact Initial Evidence
❸ Detection CVE, CWE Identifier Initial Evidence+❶❷
❹ Correlation Related CVE Identifier Initial Evidence+❶❷❸
❺ Prioritization Severity (CVSS), Exploitation (EPSS) Initial Evidence+❶❷❸
❻ Remediation Patch Tool, Code, Mitigation Advisory Initial Evidence+❶❷❸❹❺
3.2 Curriculum Pre-Training
Next, we pre-train the base LLMs (detailed in Table 9) to embed large-scale cybersecurity knowledge. We design a “curriculum” strategy that progressively introduces the corpus throughout the training process.
Curriculum Design. The “curriculum” refers to the ordered training corpus, structured by the time of threat report publication and the number of tokens in each report. Formally, let $D=\{d_{i}\}_{i=1}^{n}$ represent the corpus, where $d_{i}$ denotes individual threat reports. We first arrange $D$ based on the year and month each CVE was reported. Within each year-month grouping, the reports $\{d_{i}\}$ are further sorted by token length, creating a sequence that progresses chronologically from older to newer reports and by complexity from shorter to longer documents. For simplicity, we continue to denote the ordered corpus as $D$.
Training with Pacing. To enhance convergence and stabilize training, we introduce a pacing function Pace(,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) that governs the portion of the training corpus $D$ utilized during the $t$-th epoch. Formally:
$\textsc{Pace}(D,t)=\begin{cases}D[:|D|\cdot\frac{t}{T_{1}}]&t\leq T_{1}\\ D&T_{1}<t\leq T_{2}\\ D\cup D[\beta\cdot|D|\cdot\frac{t-T_{2}}{T-T_{2}}:]&t>T_{2}\end{cases}$ (1)
The pacing function in Eq. 1 has three stages: (a) Linear Start – In the early epochs ($t\leq T_{1}$), a progressively increasing portion of $D$ is selected for training. (b) Plateau – In the mid stage ($T_{1}<t\leq T_{2}$), as the LLM $\theta$ acquires foundational cyber threat knowledge, the entire corpus $D$ is typically used for training. (c) Reinforced End – In the final epochs ($t>T_{2}$), the corpus $D$ is augmented with the most recent threat data to reinforce training of emerging cyber threats. This augmentation is balanced by a factor $\beta$. Table 9 shows values of $T$, $T_{1}$, and $T_{2}$.
Pre-Training Objective. Identical to the pre-training objective of Llama [92], we employ Causal Language Modeling (CLM), an autoregressive training objective where the LLM predicts the next token in a sequence based on preceding tokens. Let $x_{j}$ denote a token in a threat report $d$ with $L_{d}$ tokens, we have $d=\{x_{j}\}_{j=1}^{L_{d}}$. The CLM objective is formulated as:
$\ell_{\texttt{CLM}}(\theta,D)=-\mathbb{E}_{d\sim\mathcal{D}}\left[\sum_{j=1}^{L_{d}}\log p_{\theta}(x_{j}|x_{<j})\right]$ (2)
Here, the model $\theta$ learns to maximize the likelihood of the next token $x_{j}$ given the preceding context $x_{<j}$, across all tokens in a threat report $d$. This objective enables the model to effectively internalize and represent the knowledge included in the cyber threat corpus $D$.
4 Fine-Tuning, Inference, and Customization
This section details the development of CyLens after pre-training, focusing on task-oriented dataset construction for fine-tuning (section 4.1), modularized inference (section 4.2), and organizational customization (section 4.3).
4.1 Fine-Tuning with Cascading Reasoning
Data Construction. CTI analysis follows a cascading process, which starts with an initial set of evidence and progressively gathers additional information to refine the analysis [64, 81]. To equip LLMs with cascading analytical capabilities, we design the fine-tuning dataset to reflect progressive CTI activities: $D=\{E_{0}\cup(Q_{i},A_{i})_{i=1}^{n}\}$ where $E_{0}$ represents the initial evidence, and each query-answer pair $(Q_{i},A_{i})$ corresponds to a step in the cascading reasoning process:
$(A_{1}|E_{0},Q_{1})\rightarrow(A_{2}|E_{0},A_{1},Q_{2})\rightarrow...\rightarrow(A_{n}|E_{0},A_{1},...,A_{n-1},Q_{n})$
Example 1
Below is a cascading reasoning example.
[…]
Figure 3: An illustration of modularization in CyLens, using the same threat report as described in Example 1.
Table 1 presents the analytical targets for CTI tasks ❶-❻, where each task builds upon the preceding ones. As illustrated in Example 1, identifying CVEs requires first pinpointing threat actors, which subsequently guides the prioritization.
To construct the fine-tuning dataset, we aggregate references from MITRE-CVE, NVD, and ATT&CK. For each reference, we extract available evidence as listed in Table 1 and utilize descriptive reports as the initial evidence set $E_{0}$. We then generate $(Q_{i},A_{i})$ pairs following the format exemplified in Example 1. It is important to note that not all constructed data samples encompass the full pipeline of tasks ❶-❻. Detailed dataset statistics are provided in Appendix B.
Fine-Tuning. To enable CyLens with cascading reasoning capabilities, we fine-tune it using the dataset $D$ described above. The fine-tuning process follows a standard instruction-tuning objective [106, 44], where the model is trained to generate a sequence of answers conditioned on the preceding context. Let $\theta$ represent the parameters of CyLens. The objective function is defined as:
$\begin{split}\mathcal{L}(\theta)=-\mathbb{E}_{(E_{0},Q_{1\sim n},A_{1\sim n})\in D}\ell(E_{0},Q_{1\sim n},A_{1\sim n},\theta)\\ \text{s.t.}\quad\ell=\sum_{i=1}^{n}\log P_{\theta}(A_{i}|E_{0},A_{1},\dots,A_{i-1},Q_{i})\end{split}$ (3)
where $P_{\theta}(A_{i}|E_{0},A_{1},\dots,A_{i-1},Q_{i})$ represents the probability assigned by the model to the correct response $A_{i}$ given the accumulated context up to step $i$.
Table 2: Evaluation results for ❶ Attribution tasks, where "Hist."–historical threats, "0-Day"–zero-day threats. All metrics are scaled to 0-100 (%) for consistency. The boldface and underline highlight the best and second-best results, respectively.
Threat Actor TTP Exploitation Path
Model Accuracy BERT Score Precision Recall F1 Score IoU Accuracy BERT Score
Hist. 0-Day Hist. 0-Day Hist. 0-Day Hist. 0-Day Hist. 0-Day Hist. 0-Day Hist. 0-Day Hist. 0-Day
SEvenLLM 65.94 61.17 81.25 79.83 73.44 68.62 76.85 71.53 75.11 70.04 60.14 53.90 73.56 78.42 82.19 80.65
Lily-Cyber 63.62 54.58 78.25 75.27 64.71 55.32 70.17 64.62 59.65 67.28 42.50 50.69 76.25 70.38 82.16 82.43
ChatGPT-4o 48.33 40.31 76.02 73.63 74.46 77.42 81.53 80.52 77.83 78.94 63.71 65.21 67.76 69.74 76.15 79.90
ChatGPT-o1 57.36 55.34 77.93 78.69 81.26 72.49 83.18 82.90 82.21 77.35 69.79 63.06 78.82 72.78 77.22 78.06
Gemini-Pro 61.06 57.43 81.24 78.63 78.11 79.36 81.19 82.24 79.62 80.77 66.14 67.75 62.82 69.21 75.84 74.43
Llama-405B 34.85 24.62 74.20 71.35 72.46 60.29 78.93 71.08 75.56 65.24 60.72 48.41 41.64 25.72 74.54 71.58
DeepSeek-R1 57.85 55.26 80.39 79.92 73.24 71.51 76.43 79.85 74.80 75.45 59.75 60.58 65.71 63.04 78.27 77.91
Claude-Opus 30.12 22.64 73.51 72.96 69.82 54.84 77.95 75.26 73.66 63.45 58.30 46.46 58.91 52.64 76.46 72.83
CyLens-1B 82.81 79.52 83.84 81.72 86.15 83.22 85.72 87.44 85.93 85.28 75.34 74.33 85.28 80.61 84.62 81.55
CyLens-8B 87.58 83.74 86.72 85.39 90.67 88.70 93.75 92.74 92.18 90.68 85.50 82.94 95.71 87.79 88.45 85.21
CyLens-70B 96.53 88.29 87.93 84.52 89.63 87.57 88.26 85.46 88.94 86.50 80.08 76.21 91.53 86.06 87.32 87.03
4.2 Modularized Inference
At each reasoning step $(E_{0},A_{1},\dots,A_{i-1},Q_{i})\rightarrow A_{i}$ during inference, we employ various NLP techniques to modularize the analytical process. Table 1 summarizes the NLP techniques used for each task. These techniques are implemented through the following approaches:
Topic Modeling. We implement topic modeling to interpret users’ intentions regarding the tasks they aim to perform. Specifically, we use BERTopic, a transformer-based package that eliminates the need for manually setting hyperparameters, as required in conventional methods [7, 50]. Given a threat report $E_{0}$, we implement a two-step approach: first, based on CTI tasks ❶-❻, and then, based on specific analytical targets, as outlined in Table 1.
Example 2
Topic modeling (Figure 3-(b)).
Using the same example as in 1, the topic modeling step first determines the task as threat attribution based on $Q_{1}$. It then further specifies the analytical target as the threat actor within the threat attribution task.
Named Entity Recognition (NER) and Relation Extraction. Next, we extract entities relevant to the identified topics while simultaneously identifying relationships among them. We employ the multi-agent generation approach described in section 3.1, utilizing few-shot prompting, where task-specific entities and their relationships are included as demonstrations within the prompt. This approach facilitates a structured understanding of cybersecurity threats, enhancing the accuracy and contextual relevance of extracted threat intelligence.
Retrieval Augmented Generation (RAG). We implement RAG to retrieve relevant web documents through web screening to enrich threat reports. This is especially useful for zero-
Example 3
NER and relation extraction (Figure 3-(c)).
Using Example 1, the NER identifies key entities "APT28" and "C2 communication" within the threat report. Simultaneously, relation extraction determines the "is associated with" relationship between these entities.
day threat hunting, where web-scale CTI intelligence—such as reported threat groups, affected products, and potential patches—is continuously updated. Then, we prompt GPT-o1 to rank the retrieved documents based on their relevance to the threat report $E_{0}$. To control latency, we implement a cache to store retrieved documents. Caching documents are indexed by topic modeling results (i.e., CTI tasks and analytical targets) and specific entities (e.g., APT28 or C2 communication).
Reasoning and Summarization. We implement reasoning and summarization through prompting in CyLens, where reasoning is applied when deeper analysis or multi-step thinking is required. For example, reasoning is used to analyze severity level (e.g., medium) based on gathered evidence. Summarization is then applied to synthesize insights, tailoring the output to a specific CTI task by leveraging results from topic modeling, NER, relation extraction, and RAG.
Cascading Reasoning in Inference. Since CTI tasks exhibit dependencies (Table 1), we minimize latency by concurrently executing topic modeling, NER, relation extraction, and RAG across all prerequisite tasks. Reasoning and summarization then proceed sequentially, following task dependencies to generate structured, step-by-step findings and conclusions.
4.3 Organizational Customization
Although CyLens is developed using massive CTI knowledge, it also enables customization to meet specific organizational interests on particular threat categories, specific products, or assisting an entire organization. Such customization is achieved by further fine-tuning LLMs but specifying the scope of data—for example, concentrating on cyber threats related to a single organization. This targeted customization enhances the relevance of CTI in meeting organizational needs.
We demonstrate CyLens’s potential for tailored customization in section 5.6. Through experiments, we highlight its adaptability and flexibility in addressing diverse CTI challenges across different organizational settings.
5 Evaluation Results
This section systematically evaluates CyLens’s performance on CTI tasks ❶-❻ with the following settings:
Dataset. To ensure practicality, we directly assess threats from NVD [59], MITRE [14], and Exploit-DB [62], categorizing into two groups: (1) Historical – threats reported between 2015 and 2024. (2) Zero-day – newly reported threats from the past two months during our evaluation period in 2025.
For each threat, we compile source references from [14, 59] into structured reports and prompt LLMs for analysis. The prompt structure is illustrated in Figure 4. For different tasks, we modify the system prompt and questions accordingly.
To ensure fairness, no test threats were included in training. Statistics and detailed settings are provided in Appendix C.
Baseline. We primarily consider two groups of baselines: (1) State-of-the-art (SOTA) CTI agents, SevenLLM [44] and Lily-Cyber [84], specifically developed for CTI, and (2) Industry-leading general-purpose LLMs, such as ChatGPT and Gemini, which are widely used. Additionally, we investigate the impact of ablation baselines in section 6.1.
Metrics. We use different evaluation metrics depending on the specific CTI task. The definitions, strategies, and rationale for using these metrics are provided in Appendix C.2.
Table 3: Evaluation results for ❷ Contextualization tasks, where "Hist."–historical threats, "0-Day"–zero-day threats. All metrics are scaled to 0-100 (%) for consistency. The boldface and underline highlight the best and second-best results, respectively.
Affected System (w/ Version, if available) Attack Infrastructure Impact
Model Precision Recall F1 Score IoU Precision Recall F1 Score IoU BERT Score
Hist. 0-Day Hist. 0-Day Hist. 0-Day Hist. 0-Day Hist. 0-Day Hist. 0-Day Hist. 0-Day Hist. 0-Day Hist. 0-Day
SEvenLLM 24.76 13.08 18.57 6.53 21.22 8.71 11.87 4.55 43.17 40.28 53.62 45.92 47.83 42.92 31.43 27.32 75.92 77.64
Lily-Cyber 27.73 22.42 19.62 14.57 22.9 8 17.66 12.98 9.69 44.73 35.16 56.82 53.13 50.06 42.32 33.38 26.84 73.02 76.81
ChatGPT-4o 53.76 48.62 41.80 36.09 47.03 41.43 30.75 26.13 54.51 47.78 62.55 57.22 58.25 52.08 41.10 35.20 76.52 78.16
ChatGPT-o1 58.74 55.49 58.61 47.37 58.67 51.11 41.52 34.33 61.72 50.33 65.23 63.26 63.43 56.06 46.44 38.95 78.91 80.24
Gemini-Pro 35.62 33.17 41.52 34.50 38.34 33.82 23.72 20.35 42.35 34.35 31.17 22.62 35.91 27.28 21.88 15.79 74.85 72.62
Llama-405B 11.46 8.12 21.02 10.74 14.83 9.25 8.01 4.85 31.29 28.74 27.08 20.35 29.03 23.83 16.98 13.53 72.24 68.33
DeepSeek-R1 36.71 31.88 54.63 47.62 43.91 38.19 28.13 23.60 48.92 33.96 40.92 31.87 44.56 32.88 28.67 19.68 74.26 72.41
Claude-Opus 31.26 29.82 27.35 22.41 29.17 25.59 17.08 14.67 44.65 41.71 48.90 46.48 46.68 43.97 30.44 28.18 76.32 74.15
CyLens-1B 78.64 74.45 81.24 77.26 79.92 75.83 66.55 61.07 85.73 82.24 87.62 86.61 86.66 84.37 76.47 72.96 83.75 85.26
CyLens-8B 87.91 85.82 91.86 88.12 89.84 86.95 81.56 76.92 87.46 85.71 92.83 88.17 90.07 86.92 81.93 76.87 90.03 88.31
CyLens-70B 94.54 88.17 96.04 93.75 95.28 90.87 90.99 83.28 96.37 90.16 97.69 93.18 97.03 91.65 94.22 84.58 87.33 85.84
5.1 Threat Attribution
We begin by attributing the origin of cyber threats with three targets: threat actor, TTP, and exploitation path, as outlined in Table 1. Our results yield several research findings (RF):
RF1: Specialized Development is Necessary. As shown in Table 2, industry-leading LLMs, despite their broad applicability across diverse tasks, exhibit limited performance in CTI analysis. For instance, ChatGPT-4o has achieved only
SYSTEM You are a cybersecurity threat analyst specializing in cyber threat intelligence (CTI). Your task is to assess potential attack tactics, techniques, and procedures (TTPs), and provide an intelligence assessment following structured industry frameworks such as $\cdots$ USER <Report Title> "Emerging APT Group Exploiting Zero-Day Vulnerabilities in Cloud Infrastructure." <Report Excerpt> Recent intelligence reports indicate that an Advanced Persistent Threat (APT) group has been exploiting an unpatched zero-day vulnerability in cloud service environments. The group is using spear-phishing emails to gain initial access, deploying a custom malware variant named ‘CloudPhantom,’ and leveraging compromised API keys to move laterally $\cdots$ <Question> Identify the relevant TTPs used in this attack. Let’s think step-by-step, first, determine CTI tasks among $\cdots$ ASSISTANT <CTI Task><Analytical Target><Key Entities><Entity Relations><Retrieved Evidence and Source>$\cdots$<Final Answer>
Figure 4: Prompt example for TTP analysis. For both baselines and CyLens, we prompt them to perform step-by-step analysis following the NLP modules defined in Table 1. For tasks with prior dependencies (e.g., remediation), we instruct LLMs to first complete the prerequisite tasks before proceeding.
40.31% accuracy in threat actor attribution, significantly lower than CyLens-1B (79.52%) and CyLens-8B (83.74%) for the same task. This gap underscores the necessity of specialized development, as detailed in section 3 and section 4. Tailoring models specifically for CTI significantly reduces parameter requirements—CyLens-1B and 8B are far more lightweight compared to the 500B–1T parameters of GPT-4o.
RF2: Large model size is not always beneficial. Interestingly, scaling up model size does not always translate to improved CTI effectiveness. For instance, CyLens-8B achieves 93.74% recall in TTP attribution, outperforming CyLens-70B (88.26%), suggesting diminishing returns despite undergoing the same development process. This indicates that larger models do not alaways yield better performance in CTI tasks, emphasizing the importance of domain-specific optimizations over sheer model size.
RF3: Task-specific capability varies across baselines. As shown in Table 2, baseline models exhibit varying levels of effectiveness across different CTI tasks. For instance, GPT-o1 performs obviously worse in attributing threat actors compared to exploitation paths. A deeper case study reveals that baseline LLMs struggle to accurately interpret CTI-specific terminology related to threat actors. For example:
Case Study 1
When analyzing Log4Shell (a known vulnerability) with Mirai (a botnet) as the threat actor, GPT-o1 describes the attacker as an “outside attacker”, failing to precisely identify Mirai. This demonstrates a fundamental gap in attribution accuracy for complex cyber threats.
On the other hand, GPT-o1 achieves $>$78% accuracy in analyzing exploitation paths, as this task primarily involves summarizing existing evidence instead of complex reasoning.
Case Study 2
In cases where specificity is less critical, a rough description—such as "JNDI lookup and remote code fetching"—sufficiently summarizes the Log4Shell exploitation path. While a more precise description would be: "The Log4j library contacts the attacker’s malicious LDAP server, which then responds with a reference to a remote Java class file containing malicious code," the broader summary still effectively conveys the core mechanism.
This indicates that baseline models perform better in tasks with higher generation flexibility but face challenges in tasks requiring precise CTI expertise and structured analysis.
RF4: CyLens Performs consistently across analyses. Unlike baseline models, CyLens maintains consistent performance across different analytical targets, highlighting the importance of incorporating comprehensive expertise during development. This balanced capability ensures that CyLens does not exhibit biased performance or knowledge gaps throughout the CTI lifecycle. The same conclusion is further reinforced by subsequent evaluations.
5.2 Threat Contextualization
We continue evaluating CyLens in contextualizing affected systems, attack infrastructure (e.g., domains and servers), and impact. Besides the similar findings as outlined in RF1-RF4, our evaluations derive complementary research findings:
RF5: Zero-day threat hunting is not always harder. Both Table 2 and 3 reveal that analyzing zero-day threats does not always result in worse performance compared to historical threats. For instance, when assessing the impact of cyber threats, LLMs such as CyLens-1B and SevenLLM sometimes achieve slightly better relevance in their generated analyses. We attribute this to the nature of specific CTI tasks: For impact analysis, the contextualized understanding of previously known threats allows LLMs to provide more insightful assessments, particularly when certain zero-day threats resemble threats they have already encountered. However, for tasks such as identifying affected systems, attack infrastructures, or threat actors, zero-day threats often introduce novel and previously unseen products or malware, making analysis more challenging compared to historical threats.
RF6: Larger LLMs tend to be wordy. Through detailed case study, we observe that larger LLMs tend to generate verbose responses, leading to lower text similarity as measured by BERT score. For example:
Case Study 3
CyLens-8B summarizes an impact analysis concisely: "The March 2022 cyberattack by the Lazarus Group resulted in the theft of \$620 million from the Ronin Network, significantly impacting the cryptocurrency market and…" This closely aligns with the reference [101], maintaining factual accuracy while remaining succinct. In contrast, CyLens-70B generates a more verbose response: "In March 2022, the notorious Lazarus Group orchestrated a sophisticated cyber heist, infiltrating the Ronin Network’s defenses to abscond with a staggering \$620 million in cryptocurrency. This audacious act not only sent shockwaves through the digital currency markets, causing a temporary plunge in asset values, but also undermined investor confidence…" The latter output is overly wordy and includes sentiment-heavy language (e.g., "notorious," "sophisticated," "staggering," "audacious"), making it less aligned with the original reference.
This finding aligns with RF2, reinforcing that larger models are not always more effective for CTI tasks. However, to balance overall performance across all CTI tasks, a systematic evaluation is necessary to determine the optimal model scale.
RF7: CTI agents should align with real-world practice. Despite the strong performance of SOTA agents (SevenLLM and Lily-Cyber) on certain tasks (e.g., analyzing TTPs), they struggle with others, such as attack infrastructure analysis. We attribute this to a misalignment between their training focus and CTI applications. For instance, SevenLLM is trained with a broad emphasis on NLP tasks (e.g., NER), while Lily-Cyber is developed using hand-crafted cybersecurity and hacking-related data pairs [84]. However, these training paradigms often conflate TTPs and threat actors, rather than treating them as distinct analytical components. This lack of granularity can hinder precise CTI reasoning, particularly for tasks that require structured analysis of attack properties.
Although CTI lacks a universally standardized taxonomy, structuring tasks in alignment with real-world CTI databases (e.g., MITRE-CVE and NVD) provides a more practical and reliable foundation for cyber threat analysis. This highlights the necessity of LLMs specifically designed for CTI, rather than relying on models with indirect exposure to cybersecurity knowledge or misaligned customization that does not fully reflect real-world CTI practices.
Table 4: Evaluating ❺ Prioritization performance using CVSS 3.x metrics [60]: AV–Attack Vector, AC–Attack Complexity, PR–Privileges Required, UI–User Interaction, S–Scope, C–Confidentiality Impact, I–Integrity Impact, A–Availability Impact, Base–CVSS Base Severity. We report the accuracy of correctly categorizing each metric (e.g., Attack Vector: Network). The boldface and underline highlight the best and second-best results, respectively.
Model Historical Threat Zero-Day Threat
AV AC PR UI S C I A Base AV AC PR UI S C I A Base
SL 22.51 82.64 35.04 54.21 56.87 30.14 21.62 54.67 32.75 14.76 76.85 24.46 32.27 46.83 33.70 35.62 51.84 27.58
LY 54.75 65.83 22.12 76.91 53.56 43.17 37.85 42.73 43.56 47.52 64.31 19.75 59.17 58.75 37.28 33.54 45.19 36.61
4O 76.85 92.91 61.84 80.78 66.51 27.05 20.43 34.21 58.96 77.81 84.55 56.78 62.35 68.14 22.56 31.05 28.49 46.51
O1 82.51 93.36 70.57 88.21 72.53 51.69 24.22 46.73 69.25 86.83 92.74 63.17 77.14 76.83 45.30 47.26 54.03 63.71
GM 82.07 90.71 46.42 73.72 61.24 44.75 39.80 38.19 52.51 67.38 83.26 31.07 56.34 55.62 31.27 36.56 42.93 39.81
LA 37.18 56.25 67.24 44.21 48.67 22.13 18.69 29.32 24.36 12.61 57.63 54.53 39.81 28.75 10.33 17.72 11.36 17.45
DS 64.52 87.21 54.52 87.36 69.54 46.27 44.18 40.21 62.64 58.82 82.01 46.42 82.86 51.35 35.26 40.96 38.74 53.26
CD 53.76 66.25 63.87 56.32 62.71 35.06 41.28 37.93 49.07 52.16 55.86 66.46 49.39 57.24 30.04 22.75 28.39 36.88
C1 86.26 93.32 88.27 94.55 95.51 81.10 79.63 84.23 88.17 84.71 90.83 86.06 96.64 90.32 76.64 73.59 81.62 83.44
C8 94.74 98.85 95.53 99.73 97.63 93.39 90.33 95.49 96.86 93.27 99.31 94.42 97.65 94.52 85.43 88.47 86.54 91.57
C70 94.33 100.0 97.56 97.46 98.15 94.74 92.26 95.83 97.57 92.69 100.0 97.71 95.48 93.26 90.97 94.53 92.66 93.16
5.3 Detection and Correlation
We now evaluate CyLens and baselines on ❸ detection and ❹ correlation tasks. In detection, the objective is to evaluate how effectively LLMs can identify relevant threat identifiers (CVEs and CWEs) from a given threat report. This step is crucial for enabling subsequent CTI activities, such as threat correlation and remediation. In correlation, we extend the evaluation by examining how LLMs associate other threats (CVEs) using different strategies: (i) Direct correlation–directly correlate additional CVEs. (ii) CVEs as intermediates–first detect CVEs from threat reports (from ❸) and then use them for correlation. (iii) CWEs as intermediates–first identifying CWEs (i.e., weakness categories, from ❸) and using them for correlations. The latter two strategies leverage intermediate evidence (CVEs or CWEs identified in the detection step) to uncover the threat network.
Our evaluations yield the following new research findings:
Figure 5: Evaluation results on ❸ Detection tasks using historical threats. The presentation order of LLM names (top-down) follows the same sequence as in Tables 2 (top-down).
RF8: CyLens demonstrates strong detection relevance. Figure 5-(a)(b) shows the hit ratio for the top 10 detected threat identifiers (CVEs, CWEs). Identifying CVEs is more challenging due to their vast number ($>$200k) compared to CWEs ($\sim$900). However, we also observe that CyLens  is able to detect more relevant CVEs such as those that belong to the same category or affect the same products. For example:
Case Study 4
Given the threat report: "…Apache HTTP Server allows remote attackers to execute arbitrary code. The issue arises from improper handling of user-supplied input in HTTP request headers…" corresponding to CVE-2021-42013. LLMs present different detections: GPT-o1 and DeepSeek-R1 incorrectly output CVE-2020-0609 and CVE-2022-2296, both of which are unrelated to Remote Code Execution or Apache. CyLens-70B gives CVE-2023-25690, a highly relevant HTTP Request Smuggling attack on the Apache Server.
From the above case study, although CyLens-70B does not match the exact CVE identifier, it effectively aligns with the evidence provided in the threat report and demonstrate a strong threat detection capability. This effectiveness is further supported by the results shown in Figure 5-(c)(d).
RF9: CyLens demonstrates robust threat correlation. Table 5 presents the hit ratio for the top 10 correlated threats (CVEs) as reported in MITRE-CVE [14]. We evaluate three correlation strategies: (i) Direct correlation, (ii) CVEs as intermediates, and (iii) CWEs as intermediates, as described above. We have corresponding observations:
Table 5: Evaluation results (HIT@10) on ❹ Correlation task using historical threats. The order of LLM names (left-to-right) follows the same sequence as in Tables 2 (top-down).
SL LY 4O O1 GM LA DS CD C1 C8 C70
Direct 0.6 0.0 3.5 8.1 2.2 0.0 0.0 0.4 52.7 74.3 77.5
w/ CVE 7.3 4.5 11.6 25.4 6.7 0.0 4.3 12.5 60.4 79.0 81.2
w/ CWE 0.0 0.0 0.0 0.8 0.0 0.0 0.0 0.0 41.7 70.8 71.6
    • (i)
      While baseline models largely struggle to correlate CVEs, CyLens demonstrates significant effectiveness (e.g., 77.5% Hit@10 with CyLens-70B).
    • (ii)
      Incorporating CVEs as intermediate evidence improves correlation performance. However, baseline models tend to propagate erroneous CVEs from the prior detection step, which misguides subsequent threat correlations.
    • (iii)
      Incorporating CWEs degrades correlation performance due to their inherent ambiguity. Since a CWE can broadly relate to many CVEs, it introduces ambiguity, impeding accurate threat correlation. Despite this challenge, CyLens remains robust (e.g., 71.6% Hit@10 with CyLens-70B) in incorporating ambiguous intermediate evidence while maintaining strong correlation effectiveness.
5.4 Threat Prioritization
Real-world cyber threats often co-occur across cyber events [77, 73], necessitating that CTI teams distinguish threat severity and prioritize subsequent patches and mitigations. Our evaluation focuses on two prioritization measurements: severity level and analysis of exploitation dynamics.
Severity Level. Threat severity can be assessed from threat reports even when CVE identifiers are not disclosed or when zero-day threats have not yet been assigned CVEs. We evaluate severity using 8 metrics defined in CVSS 3.x [60], which collectively capture a threat’s exploitability and impact. Each severity metric is treated as a classification task. For example, the Attack Vector metric has four predefined categories: Network, Adjacent Network, Local, and Physical. In cases where a threat report lacks sufficient evidence for CVSS metrics, we introduce an additional category, "N/A", for each metric.
Furthermore, we classify the base severity level from four categories: Low, Medium, High, and Critical. This classification aggregates the ratings from 8 CVSS metrics, providing an overall severity assessment aligned with CVSS standards.
RF10: CyLens maintains stable effectiveness. Table 4 shows the accuracy of all models. While baselines vary widely across CVSS metrics (e.g., GPT-o1 scores 92.91% on Attack Complexity but only 20.43% on Integrity Impact), CyLens remains consistently strong. Notably, CyLens-8B achieves $>$90% accuracy across 8 metrics for historical threats, demonstrating its reliability in handling different severity factors.
This stability comes from CyLens ’s ability to adapt to different evidence types. Each CVSS metric relies on specific context—Privileges Required (PR) depends on access control policies, while Attack Complexity (AC) is influenced by real-world constraints and attacker skill level. By effectively integrating diverse evidence sources, CyLens reduces inconsistencies in severity analyses, which aligns with RF4.
Additionally, we find that zero-day threats are not always harder to analyze, aligning with RF5. Some evidence, like some user actions in exploitations, is already known from past threats, facilitating assessment on newly emerging threats.
We also observe that larger LLMs do not always perform better, consistent with RF6. While bigger models tend to provide more detailed reasoning, this can sometimes lead to inaccurate analysis. For example, CyLens-70B does not always outperform CyLens-8B across all severity metrics.
Exploitation Dynamics. Besides severity analysis using CVSS, we also evaluate CyLens’s effectiveness in analyzing dynamic exploitation behaviors. Specifically, we use the EPSS score [29] and conduct two types of analysis: (1) pinpointing historical exploitation at an earlier timestamp and (2) forecasting future exploitation at a later timestamp.
Figure 6: With $\mathcal{T}_{0}$ as the “current” timestamp (varying among threats), we compute RMSE for historical EPSS scores over the past three months (interpolation) and future EPSS scores for the next three months (prediction) w.r.t. $\mathcal{T}_{0}$.
For test threats with more than six months of EPSS records, we set a reference timestamp $\mathcal{T}_{0}$ for each cyber threat (e.g., 01/01/2024) and collect EPSS reports from $\mathcal{T}_{0}$-3 months (e.g., 10/01/2023) to $\mathcal{T}_{0}$+3 months (e.g., 04/01/2024). However, EPSS reports between $\mathcal{T}_{0}$-3 to $\mathcal{T}_{0}$ are included in the inference prompt as prior knowledge, similar to the prompt structure illustrated in Figure 4.
We compare CyLens-1B/8B against four baseline models and compute RMSE between the interpolated (historical) or predicted (future) EPSS scores and the actual scores (scaled from 0% to 100%). To ensure accurate evaluation, we match each EPSS score to the closest available recorded score. For example, if an EPSS score is not reported exactly on 10/01/2023, we use the closest available date (e.g., 09/25/2023) from EPSS reports [14] as the ground truth.
RF11: LLMs vary in interpolation and prediction. As shown in Figure 6, LLMs perform slightly better in interpolating historical EPSS scores (lower RMSE) than in predicting future scores. However, we also observe that local LLMs, such as SevenLLM, perform obviously worse in both interpolation and prediction due to their static inherent knowledge, which lacks adaptability to evolving cyber threats. On the contrary, CyLens, despites its local deployment, performs similar to API-based LLMs (e.g., Gemini), which demonstrates the necessity of equipping LLMs with domain expertise. This enables LLMs to adapt to the dynamic nature of CTI tasks.
RF12: Contextual knowledge is essential. Beyond Figure 6, we observe that some cyber threats undergo dramatic shifts in EPSS scores due to emerging exploitation activities.
Case Study 5
For instance, CVE-2024-27956 saw a sharp increase in EPSS on 02/15/2025, jumping from 0.31% to 55.37% [19]. No LLM was able to accurately predict this trend without contextual knowledge of real-world events.
This implies the importance of not only fitting to EPSS trends but also incorporating real-world evidence in real time to enhance CTI adaptability.
5.5 Remediation
Table 6: Evaluation results on ❻ Remediation. We evaluate whether LLMs can correctly generate: (i) patching or mitigation tools – measured by accuracy, (ii) code patches – measured by syntactic similarity, (iii) patching or mitigation methods – measured by BERT Score, and (iv) reference advisories – measured by accuracy.
Model Tool Use Code Patch Methodology Advisory
Hist. 0-Day Hist. 0-Day Hist. 0-Day Hist. 0-Day
SL 87.81 79.65 81.83 76.62 82.95 82.48 56.74 43.25
LY 79.63 57.17 75.96 73.60 80.75 78.53 48.71 46.63
4O 86.62 74.51 82.70 81.05 84.42 82.51 76.83 71.75
O1 90.08 86.65 84.47 82.86 85.53 85.25 81.26 77.79
GM 76.83 53.22 83.14 82.47 84.79 80.32 66.72 43.25
LA 75.71 69.20 81.23 78.46 83.26 82.59 49.53 38.17
DS 73.70 70.21 79.27 77.51 83.64 82.55 65.43 47.92
CD 77.54 73.48 76.19 73.64 82.27 81.92 54.85 47.35
C1 92.03 88.84 85.42 83.71 87.83 84.52 88.74 82.25
C8 93.83 92.74 87.48 87.36 89.16 88.37 92.32 88.54
C70 97.62 97.23 90.66 86.39 91.20 88.04 95.34 92.26
Finally, we evaluate CyLens’s capability in generating accurate remediation strategies, focusing on four aspects: suggesting the correct tool for patching or mitigation (e.g., OpenSSL for encryption), generating available code patches (e.g., GitHub commits), proposing appropriate mitigation methodologies (e.g., hardening an OS), and identifying the correct advisory to consult (e.g., NVIDIA Security Advisory).
RF13: CyLens is effective across all remediation tasks. Consistent with RF4 and RF10, we demonstrate CyLens’s effectiveness in handling various remediation objectives as shown in Table 6, including both identification-based (suggesting tools and advisories) and generation-based (producing code patches or mitigation strategies) approaches. In contrast, baseline models, lacking CTI-specific expertise, often exhibit imbalanced performance across remediation tasks. For example, Gemini performs well in generating code patches but struggles to identify the correct advisories.
Figure 7: Evaluation results across multiple CTI tasks based on different levels of organizational interests: (a)-(d) product-level, (e)-(h) vendor-level, and (i)-(l) threat category-level. The evaluation metrics used are accuracy for threat actor attribution, F1 score for attack infrastructure contextualization, accuracy for CVSS base severity classification, and accuracy for third-party advisory suggestion. Model names are SL: SevenLLM, O1: ChatGPT-O1, GM: Gemini-Pro, C1/C8: CyLens-1B/8B.
RF14: Large-scale expertise is crucial. Effective remediation often relies heavily on existing cybersecurity practices [43]. Therefore, embedding large-scale historical remediation knowledge is crucial when developing LLMs for CTI applications. Even for zero-day threats, well-trained LLMs can leverage their inherent knowledge to derive useful insights, improving response effectiveness even in cases where no explicit remediation has been documented.
5.6 Organizational Customization
Previous experiments demonstrate CyLens’s capability in analyzing a wide range of cyber threats, providing a strong foundation for addressing specific organizational needs. Given that threats of particular interest—such as those associated with specific products, vendors, or threat categories—may lack sufficient volume for fine-tuning [58, 35], it is essential to first develop LLMs with comprehensive CTI analysis capabilities. These LLMs can then be incrementally customized to align with specific organizational needs.
We categorize organizational needs into three levels:
    • Product-Level – Focusing on cyber threats affecting specific products, a common approach among security teams in large enterprises managing product-specific CTI [21].
    • Vendor-Level – Targeting threats that impact an entire vendor (i.e., company), a perspective often adopted by small and medium-sized businesses or industry groups [67].
    • Category-Level –Concentrating on specific threat categories, which is particularly relevant to research institutions conducting specialized cybersecurity studies [49].
Settings. To cover diverse organizational needs, we collect threats associated with (1) specific products (Acrobat, Chrome, or Fedora), (2) specific vendors (openSUSE, GitLab, or Dell), and (3) specific threat categories (Overflow, DoS, or XSS). Thus, we collect a total of nine groups of threats, with each group comprising more than 1,000 threat reports. We allocate 50% of these threat reports—sourced from third-party references in the MITRE-CVE [15], NVD, and Exploit-DB [62] websites [59]—for independently fine-tuning CyLens-1B/8B, while reserving the remaining for evaluation. We compare CyLens to the SOTA model, SevenLLM, and two industry-leading LLMs, GPT-o1 and Gemini. Figure 7 presents the results across multiple CTI tasks.
RF15: CyLens is adaptable to diverse organizational needs. From Figure 7, we observe that CyLens maintains high effectiveness even with independent customization (fine-tuning) across different groups of cyber threats. For instance, CyLens-8B achieves nearly 100% accuracy in providing remediation advisories for various vendors and threat categories. This result further demonstrates its adaptability despite the diverse nature of threats across products, vendors, and categories—factors that significantly impact baseline LLMs, causing their performance to fluctuate. The finding implies CyLens ’s superior capability to generalize while maintaining high effectiveness in tailored needs.
6 Discussion
6.1 Ablation Study
We conduct ablation studies to assess the impact of different components within CyLens. Table 7 presents the results, where we independently exclude pre-training (PT), fine-tuning (FT), and four optional modules: topic modeling (TOM), named entity recognition (NER), relation extraction (REL), and retrieval-augmented generation (RAG). Notably, reasoning (REA) and summarization (SUM) cannot be excluded, as they are essential steps in text generation.
Table 7: Ablation study on CyLens-8B by selecting four representative CTI tasks: (1) attributing threat actors (measured by accuracy), (2) contextualizing attack infrastructure (measured by F1 score), (3) classifying CVSS base severity levels (measured by accuracy), and (4) suggesting advisories for remediation (measured by accuracy).
Ablation Threat Actor Attack Infra CVSS Base Advisory
Hist. 0-Day Hist. 0-Day Hist. 0-Day Hist. 0-Day
w/o PT 66.22 62.31 70.05 61.44 53.62 45.90 57.25 54.36
w/o FT 82.42 70.15 86.47 78.31 86.48 61.75 82.56 77.23
w/o TOM 84.23 81.17 88.46 82.24 95.12 88.82 87.23 85.44
w/o NER 71.12 70.19 82.83 72.63 94.52 91.42 82.53 80.65
w/o REL 87.21 83.09 88.18 86.21 95.67 91.25 83.12 82.72
w/o RAG 84.17 80.26 86.53 82.71 95.23 89.64 85.14 83.42
We observe that pre-training has the most significant impact, as it embeds massive CTI knowledgeinto the model. While fine-tuning is more lightweight, its influence remains non-negligible in refining CyLens ’s CTI capabilities. Additionally, the impact of different modules varies across tasks. For instance, classifying CVSS severity levels does not heavily depend on NER and REL, so removing these modules has minimal effect. However, tasks such as threat actor attribution or advisory suggestion rely on accurately identifying key objects (e.g., malware or third-party vendors), making NER and REL critical for performance. These findings reinforce the necessity of incorporating all proposed techniques to achieve an effective and well-rounded system design for CTI applications.
6.2 Adapting to Computational Constraints
Similar to Meta’s Llama series, which offers models at different scales [56, 57, 92], CyLens is available in three sizes: 1B, 8B, and 70B, allowing it to accommodate varying computational constraints. While CyLens-1B is less effective compared to the 8B and 70B versions, its lightweight design is crucial for resource-limited deployments.
In scenarios where computational resources are highly constrained, having a functional LLM is often more practical than striving for state-of-the-art performance. Notably, despite its smaller size, CyLens-1B still outperforms industry-leading LLMs in most CTI tasks, reinforcing its value as an efficient yet capable model for cybersecurity applications.
6.3 Privacy-Preserving Deployment
A key advantage of customizing local LLMs is enhanced privacy preservation. Unlike API-based services such as OpenAI’s GPT series, which inherently risk exposing user intentions and confidential data, local models ensure that sensitive information remains under organizational control. API-based solutions may not only transmit proprietary data to external servers but could also use it for model training and further development, raising concerns about data security and regulatory compliance [65, 66, 80, 99, 113].
By developing localized, task-specific LLMs, organizations gain full control over model behavior, customization, and data security. Such privacy-preserving deployment is particularly critical in high-stakes fields like CTI, where threat intelligence, incident reports, and security strategies must remain confidential to prevent adversarial exploitation and ensure compliance with data protection regulations.
6.4 Beyond Specific Needs
Real-world cyber threats are rarely confined to specific products, vendors, or categories [77]. Instead, they manifest within a complex and interconnected ecosystem where vulnerabilities, exploits, and attack methodologies transcend traditional boundaries. This interdependence underscores the necessity of developing CTI solutions that go beyond addressing specific organizational needs (e.g., tailored threat intelligence) and are also capable of capturing the broader, multifaceted threat landscape.
7 Conclusion
This paper introduced CyLens, a LLM-powered copilot for cyber threat intelligence. CyLens streamlines the entire threat management lifecycle, encompassing tasks from threat attribution to remediation. It assimilates knowledge from a vast number of threat reports and utilizes a retrieval module for real-time information, enabling it to adeptly handle both historical and zero-day threats. Rigorous evaluations have demonstrated that CyLens consistently outperforms industry-leading LLMs and SOTA agents across multiple CTI tasks. Its adaptability shines through its customization capabilities, which can be tailored to specific threat categories or organizational requirements. Additionally, the release of its training datasets fosters further research in the field.
Appendix A Pre-Training: Complementary Information
Table 8: The data sources we utilize and the corresponding evidence metadata collected from each database.
CTI Task Analysis Target Source Database
CVE NVD CWE ATT&CK CAPEC Exploit DB Third-Party Ref.
Threat Actor
TTPs
Exploit Path
Affected System
Attack Infra
Attack Impact
CVE Identifier
CWE Identifier
CVE Identifier
CVSS Metrics
EPSS Records
Tool Use
Code Patch
Methodology
Advisory
Table 9: Base LLM assignment and settings of CyLens’s development, including pre-training and fine-tuning. “Cp” – Corpus; “Tk” – token.
CyLens Base LLM Pre-Training
# Cp # Tk $T$ $T_{1}$ $T_{2}$ $\beta$
CyLens-1B Llama-3.2-1B 28K 0.27B 4 3 3 1.0
CyLens-8B Llama-3.1-8B 115K 1.1B 10 5 5 1.0
CyLens-70B Llama-3.3-70B 271K 2.5B 20 10 10 1.0
A.1 Data Source and Metadata Collection
Complementary to section 3.1, Table 8 provides a detailed overview of the cyber threat databases we utilize, along with the corresponding metadata collected for various CTI tasks (❶-❻) and analytical targets.
MITRE-CVE and NVD. Our collected data is primarily centered around CVEs, leveraging the MITRE-CVE [14] and NVD [59] databases, which serve as well-structured repositories of publicly disclosed vulnerabilities. These databases provide foundational information, including CVE identifiers, descriptions, severity scores, affected products, and references to external sources.
Exploit-DB. To further enrich our dataset and enhance contextual understanding, we incorporate additional cyber threat intelligence (CTI) sources, including Exploit-DB [62]. Exploit-DB provides reports on security advisories and threat intelligence insights, complementing the structured data from MITRE-CVE and NVD. Notably, Exploit-DB is also CVE-centric; however, unlike MITRE-CVE and NVD, it lacks structured connections to additional evidence. Therefore, our data collection primarily focuses on extracting raw threat descriptions and attack vector classifications (e.g., "Local") to supplement our dataset with additional contextual details.
Other MITRE Databases. Beyond these primary databases, we integrate metadata from additional sources to enrich our dataset with diverse perspectives on cyber threats. Specifically, we leverage CWE [16] to systematically categorize software weaknesses, providing a structured foundation for understanding vulnerability root causes. CAPEC [17] offers a comprehensive taxonomy of attack patterns, enabling us to analyze how adversaries exploit known weaknesses in real-world scenarios. MITRE ATT&CK [18] serves as a crucial resource for mapping adversary tactics, techniques, and procedures (TTPs), facilitating in-depth attribution and behavioral analysis.
Third-Party References. Beyond structured frameworks, we incorporate third-party references to capture real-time threat intelligence and contextualized security insights. Among them, GitHub commits to serve as a valuable source for tracking vulnerability disclosures, exploit development, and security patch discussions. Apache security reports [5] provide detailed documentation on vulnerabilities and remediation efforts specific to widely used web and enterprise software. Meanwhile, IBM X-Force threat intelligence [39] contributes curated threat reports, real-world attack observations, and insights into evolving threat landscapes.
The collected metadata, detailed in Table 8, enhances the comprehensiveness of our dataset by offering multifaceted insights into vulnerability exploitation patterns, adversarial attack techniques, and effective mitigation strategies. By integrating these diverse sources, we enable a richer and more contextualized analysis of cyber threats, supporting improved detection, attribution, and defense mechanisms.
A.2 Threat Corpus Generation Details
Complementary to section 3.1, we detail the prompts used for generating and refining the threat corpus, specifically in Prompt 1 and 2. During output sampling, we adjust the temperature and top-p parameters within a range of 0.2 to 1.0 to balance creativity and determinism. An example of a generated threat report is provided in Example 4. Additionally, we present the statistical overview of the corpus used for pre-training in Table 9.
Prompt 1
Below is the prompt template used for threat corpus generation.
[…]
Prompt 2
Below is the prompt used for threat corpus revision.
Objective: You are an expert cybersecurity analyst and technical writer. Your task is to revise the provided cyber threat report by improving its layout and structure while maintaining all critical details. The goal is to $\cdots$ Instructions: Reorganize Content – Adjust the structure to make the report more logical and easier to follow. Consider using sections such as $\cdots$ Improve Readability – Use concise paragraphs, bullet points, and tables where appropriate. Ensure technical jargon is well-explained. Maintain Accuracy – Preserve all details, technical data, and intelligence without altering their meaning. $\cdots$ Report to Revise: [Insert Cyber Threat Report Here] Output:
Example 4
Below is an example of threat report.
Title: Security Advisory: Cleartext Storage Vulnerability in ChatGPT macOS Application Executive Summary: A security vulnerability has been identified in versions of the ChatGPT application for macOS released before July 5, 2024. The application was found to store user conversations in cleartext, making them accessible to other applications on the system. Additionally, the app operated outside of the macOS sandbox, $\cdots$ Vulnerability Details: CVE Identifier: CVE-2024-40594 Description: The ChatGPT macOS application prior to version 2024-07-05 stored user conversations in an unencrypted format in a location accessible by other applications. Furthermore, the application opted out of the macOS sandbox, which is designed to limit the app’s access to system resources and user data $\cdots$ Affected Products: $\cdots$ Impact: The cleartext storage of conversations poses a privacy risk, as sensitive information could be accessed by other applications or users with sufficient privileges on the same system. The lack of sandboxing further exacerbates the issue by allowing the application broader access to system resources than necessary, potentially exposing additional data $\cdots$ Mitigation: Users are advised to update the $\cdots$
Appendix B Fine-Tuning: Complementary Information
Complementary to section 4.1, we present the statistics of the generated fine-tuning dataset in Table 10. Notably, when generating each data instance using a cascading reasoning layout, not all analytical targets are included due to variations in their availability.
Table 10: Statistics of the fine-tuning dataset with a cascading reasoning layout, where each analytical target appears in varying volumes of data. Additionally, different scales of fine-tuning data are used for different CyLens models to align with their respective capacities and optimization needs.
CTI Task Analytical Target # Sample # Fine-Tuning Samples
1B 8B 70B
Threat Actor 37.6K 0.8K 11.5K 37.6K
TTPs 35.5K
Exploit Path 21.8K
Affected System 37.6K
Attack Infra 28.3K
Attack Impact 17.6K
CVE Identifier 37.6K
CWE Identifier 37.6K
CVE Identifier 22.9K
CVSS Metrics 37.6K
EPSS Records 33.1K
Tool Use 15.3K
Code Patch 8.2K
Methodology 32.4K
Advisory 27.5K
Furthermore, CyLens-1B, CyLens-8B, and CyLens-70B utilize different volumes of fine-tuning data, reflecting their respective model capacities and optimization requirements. For fine-tuning CyLens-1B, we select 0.8K samples that contain all analytical targets. In contrast, for CyLens-8B and CyLens-70B, we use a progressively larger dataset, with CyLens-8B utilizing a partial subset and CyLens-70B leveraging the full dataset, despite the potential absence of certain analytical targets in some instances.
Appendix C Evaluation: Complementary Information
C.1 Evaluation Data
Rationale. Evaluating CTI using both historical cyber threats and zero-day threats provides a comprehensive assessment of their robustness and adaptability. Historical threats serve as a fundamental benchmark, allowing for the validation of defensive mechanisms against well-documented attack patterns, known TTPs. This retrospective analysis ensures that security measures remain effective against previously encountered threats and can detect variations of known exploits. On the other hand, zero-day threats introduce an element of unpredictability, testing the system’s ability to detect and respond to novel attack vectors without predefined signatures or heuristics.
By incorporating both types of threats, evaluations can measure not only the efficiency of existing threat detection models but also the adaptability of security solutions in dynamic and evolving threat landscapes. This dual approach is essential for developing resilient cybersecurity frameworks that balance historical insights with forward-looking threat intelligence.
Table 11: Statistics of evaluation dataset
Historical Zero-Day
# Threats 10,000 5,526
Table 11 presents the volume of our evaluation datasets. We ensure that no threats included in the evaluation appear in the training data.
C.2 Evaluation Metrics
We define the evaluation metrics used for each analytical target, along with the strategies and rationale behind their selection.
    • Threat Actor: Since threat actors are typically unique in each threat report, we use Accuracy to assess whether the threat actor generated by LLMs matches the reference. Additionally, we employ BERT Score to measure the textual similarity between the generated explanation and the reference for threat actor attribution. BERT Score is implemented using the Hugging Face library [^2] https://huggingface.co/spaces/evaluate-metric/bertscore.
    • TTPs: As TTPs are typically presented as a list, we adopt four metrics:
          Precision – Measures the proportion of correctly identified TTPs among those predicted by the LLM.
          Recall – Measures the proportion of actual TTPs that were successfully identified.
          F1 Score – Provides a balanced measure of precision and recall.
          Intersection over Union (IoU) – Quantifies the overlap between the predicted and reference TTP sets.
    • Exploitation Path: Since exploitation paths are usually descriptive observations, we use BERT Score to measure text similarity between the generated and reference descriptions.
    • Affected System & Attack Infrastructure: Similar to TTPs, both affected systems and attack infrastructures are structured as itemized lists. We therefore apply Precision, Recall, F1 Score, and IoU to assess the overlap between the LLM-generated results and the reference lists.
    • Impact: As impact analysis consists of descriptive statements, we use BERT Score to evaluate textual similarity between generated and reference statements.
    • CVE and CWE Detection & Correlation: Given the large-scale CVE candidates and hundreds of CWEs, the detection and correlation tasks prompt LLMs to generate the top-10 most relevant identifiers (CVE or CWE). We use Hit@10 to measure whether at least one correct identifier is present in the top-10 results.
    • Correct Product & Threat Category: To assess whether the generated CVEs correspond to the correct affected product or fall within the correct threat category, we compute Accuracy as the evaluation metric.
    • CVSS Metrics: Since CVSS metrics have a limited number of categorical values (e.g., base severity levels), we use classification accuracy to evaluate the correctness of the predicted severity levels.
    • EPSS Trend Prediction: Since the EPSS score is a numerical probability (e.g., a 5.5% chance of exploitation within a given timeframe), we calculate the Root Mean Square Error (RMSE) to measure the deviation between the LLM-predicted score and the ground truth.
    • Tool Use & Advisory Identification: Both of these tasks involve selecting a single correct item, so we evaluate performance using Accuracy to measure whether the LLM can identify the correct tool or advisory.
    • Code Patch & Mitigation: Since both tasks involve generating textual recommendations, we use BERT Score to assess the similarity between the LLM-generated output and the reference.

### v2

@@
-Cyber Defense Reinvented: Large Language Models as Threat Intelligence Copilots
-Abstract
+CyLens: Towards Reinventing Cyber Threat Intelligence in the Paradigm of Agentic Large Language Models
+Abstract.
 The exponential growth of cyber threat knowledge, exemplified by the expansion of databases such as MITRE-CVE and NVD, poses significant challenges for cyber threat analysis.
-Security professionals are increasingly burdened by the sheer volume and complexity of information, creating an urgent need for effective tools to navigate, synthesize, and act on large-scale data to counter evolving threats proactively.
-However, conventional threat intelligence tools often fail to scale with the dynamic nature of this data and lack the adaptability to support diverse threat intelligence tasks.
-In this work, we introduce CyLens, a cyber threat intelligence copilot powered by large language models (LLMs).
-CyLens is designed to assist security professionals throughout the entire threat management lifecycle, supporting threat attribution, contextualization, detection, correlation, prioritization, and remediation.
+Security professionals are increasingly burdened by the sheer volume and complexity of information, creating an urgent need for effective tools to synthesize massive knowledge and counter evolving threats proactively.
+However, conventional threat intelligence often struggle to scale with the rapidly expanding threat landscape and lack the adaptability required to support diverse threat intelligence tasks.
+In this work, we introduce CyLens, a cyber threat intelligence system powered by large language models (LLMs).
+CyLens is designed to assist security professionals throughout the entire threat analysis lifecycle, supporting threat attribution, contextualization, correlation, prioritization, and remediation.
 To ensure domain expertise, CyLens integrates knowledge from 271,570 threat reports into its model parameters and incorporates six specialized NLP modules to enhance reasoning capabilities.
@@
 By detailing its design, development, and evaluation, this work provides a blueprint for leveraging LLMs to address complex, data-intensive cybersecurity challenges.
-1 Introduction
+Cyber Threat Intelligence, Large Language Models, Threat Hunting
+* Xiaoqun Liu and Jiacheng Liang contributed equally to this work.
+copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/2018/06ccs: Security and privacy Systems security
+1. Introduction
 We are living in an era of rapid digital transformation, where technological advancements are closely linked to the growing prevalence of cyber threats.
-In recent years, the landscape of cyber threats has evolved dramatically, marked by a 25% annual increase in reported CVEs [40].
-The increase in CVEs can be attributed to several factors, especially the growing complexity of IT systems [76], the widespread adoption of open-source software [22, 79], and the accelerated pace of development cycles [83].
+In recent years, the landscape of cyber threats has evolved dramatically, marked by a 25% annual increase in reported Common Vulnerabilities and Exposures (CVEs) (Intelligence, 2023).
+The increase in CVEs can be attributed to several factors, especially the growing complexity of IT systems (Qualys, 2024), the widespread adoption of open-source software (Dam and Neumaier, 2023; Reading, 2024), and the accelerated pace of development cycles (Security, 2024).
 Together, these factors facilitate the exploitation of cyber threats in modern software ecosystems.
-To enable efficiently analyze and counteract these threats, Cyber Threat Intelligence (CTI) [23] has emerged as a crucial strategy.
-CTI provides actionable insights derived from the analysis of threat data and empowers security teams to anticipate, identify, and mitigate cyber risks effectively [89, 110, 8].
-However, CTI faces critical challenges in scalability and adaptability.
-First, the overwhelming volume of threat data creates significant bottlenecks [1].
+Figure 1. (Left) An agentic LLM execution on threat intelligence tasks. (Right) The customization workflow to enhance the CTI proficiency of pre-trained LLMs.
+[…]
+Introduction Example.
+To efficiently analyze and counteract these threats, Cyber Threat Intelligence (CTI) (Dehghantanha et al., 2018) has emerged as a crucial strategy.
+CTI provides actionable insights derived from the analysis of threat data and empowers security teams to anticipate, identify, and mitigate cyber risks effectively (Sun et al., 2023; Bromiley, 2016).
+However, it faces two critical challenges: (1) Scalability: The overwhelming volume of threat data creates significant bottlenecks (Abu et al., 2018).
 Security teams are often inefficiently processing and correlating this massive influx of security knowledge, limiting their ability to respond to emerging cyber threats.
-Second, integrating CTI into existing cybersecurity platforms often requires extensive customization, which poses a significant barrier for small and medium-sized organizations who lack resources to deploy CTI effectively.
-To address these challenges, it is essential to propose advanced CTI solutions capable of scalably collecting and analyzing extensive cybersecurity knowledge while adaptively support customization tailored to diverse organizational needs.
-Figure 1: (a) The lifecycle of threat management.
-Each task consists of multiple analytical targets, as detailed in Table 1.
-(b) The development methodology of CyLens, as detailed in section 3 and section 4.
-Large language models (LLMs) have recently demonstrated exceptional capabilities, driven by their scalability to massive knowledge through pre-training and their adaptability for organization-specific applications via fine-tuning.
-Beyond these strengths, LLMs can be further augmented with task-oriented prompts or external modules, enabling more advanced analyses that require topic modeling [96], entity and relation recognition [97, 37], and retrieval-augmented generation (RAG) [33, 10].
-Hence, LLMs emerge as a promising solution for assisting humans in specialized fields such as bug fixing [72, 94] and fuzz testing [38, 104, 24].
-Consequently, we raise a critical question: how can LLMs be systematically developed as copilots to support CTI activities?
-This work.
-To address this question, we propose CyLens, an LLM-powered copilot system designed to enhance CTI.
-CyLens consists of several sizes—1B, 8B, and 70B—to accommodate diverse needs and computational constraints.
+(2) Adaptability: Integrating CTI into existing cybersecurity platforms often requires extensive customization, which poses a significant barrier for small and medium-sized organizations who lack resources to deploy CTI effectively.
+To address these challenges, it is essential to propose advanced CTI solutions that adaptively support customization tailored to diverse organizational needs.
+Figure 2. The lifecycle of threat intelligence tasks and running examples.
+[…]
+CTI pipeline and example.
+Large language models (LLMs) have recently demonstrated exceptional capabilities in natural language understanding and reasoning.
+Compared to conventional CTI solutions, such as rule-based systems (Ali and Peng, 2024; Sun et al., 2023) or function-frozen AI models (Montasari et al., 2021; Kumar et al., 2024; Kant and Amrita, 2023; Dutta and Kant, 2020), LLMs offer superior scalability, as they can generalize across diverse knowledge-intensive domains without the need for task-specific architectures or manually crafted features (Budnikov et al., 2024; Kunter et al., 2013; Brown, 2020).
+LLMs also exhibit strong adaptability.
+Techniques like instruction tuning and few-shot learning enable them to respond effectively to dynamic and data-scarce environments, where traditional models often struggle to keep pace (Liu et al., 2024; Paeedeh et al., 2024; Wang et al., 2021; Al-Shedivat et al., 2017).
+However, despite recent studies showing that industry-leading LLMs are not yet fully qualified for CTI analysis (Srikanth et al., 2024; Kouremetis et al., 2025; Xu et al., 2024; Shafee et al., 2024), the blueprint for systematically customizing LLMs to support scalable and adaptive CTI analysis remains largely unexplored.
+This leads us to a central question: How can LLMs be systematically developed to support diverse CTI activities?
+Our work.
+To address this question, we propose CyLens, an LLM-powered agentic system designed to enhance CTI.
+CyLens is available in multiple model sizes—1B, 8B, and 70B—to accommodate varying computational constraints.
+As illustrated in Figure 1, CyLens follows the Agentic LLM paradigm, inheriting pre-trained language understanding and reasoning capabilities, and undergoing a series of targeted customizations to improve domain-specific expertise.
 We characterize CyLens by the following aspects:
-Lifecycle-Oriented Threat Management – As illustrated in Figure 1-(a), CyLens supports six tasks that cover the entire lifecycle of threat management.
-To analyze the properties of threats, CyLens enables ❶ attribution by identifying threat actors and gathering technical evidence, such as TTPs and exploitation paths, to trace the origin of a threat.
-It also facilitates ❷ contextualization by enriching metadata—such as affected systems—to provide a more comprehensive understanding of the threat landscape.
-Next, ❸ detection involves specifying available identifiers, such as CVE [14], while ❹ correlation enables CyLens to uncover adversarial networks by identifying relationships between the current threat and others.
-CyLens then drives ❺ prioritization by quantifying the severity and predicting the likelihood of exploitation.
-Building on this analytical foundation, ❻ remediation effectively counteracts threats and limits their broader impact.
-Collective Expertise.
-As shown in Figure 1-(b), CyLens manages web-scale CTI knowledge through a combination of development methodologies.
-Our approach is motivated by the fact that cyber threats co-exist as an interdependent community, where attack vectors, infrastructure, and exploitations are deeply interconnected [46, 6, 93, 12, 103].
-Effectively addressing these complex relationships requires a vast, integrated set of expertise rather than isolated knowledge.
-To achieve this, CyLens first undergoes pre-training on a diverse set of databases—MITRE-CVE [14], CWE [16], CAPEC [17], ATT&CK [18], NVD [59], Exploit-DB [62], and correlated third-party repositories (e.g., [39, 5])—to incorporate extensive CTI knowledge from 271,570 threat reports.
-We introduce a curriculum pre-training approach, where “curriculum” data is designed to progressively embed CTI expertise into CyLens.
-Following pre-training, CyLens undergoes fine-tuning using task-oriented datasets addressing critical CTI tasks (❶-❻).
-To enhance its analytical depth, we employ a cascading reasoning process, allowing CyLens to systematically build awareness of dependencies within CTI analysis.
-Additionally, we integrate specialized design modules to reinforce CyLens ’s ability to conduct rigorous reasoning, escalating it as a human-usable CTI copilot.
-Operational Proficiency.
-Our experiments demonstrate CyLens’s effectiveness across all above tasks under two categories of cyber threats: (i) Historical Threats—previously recorded threats that security analysts may not be fully aware of, and (ii) Zero-Day Threats—newly emerging threats that have not yet been documented.
-Across both threat categories, CyLens consistently outperforms SOTA models and industry-leading LLMs (e.g., ChatGPT and Gemini), underscoring its versatility in handling diverse cybersecurity challenges while excelling in proactive threat hunting within the ever-evolving cyberattack landscape.
-Adaptive to Customization.
-CyLens also demonstrates adaptability in addressing organization-specific threats related to particular products, vendors, or threat groups.
-As shown in our evaluations in section 5.6, CyLens consistently delivers superior and balanced performance across diverse organizational requirements.
-This adaptability is particularly valuable when an organization’s threat-related data volume is insufficient, enabling effective CTI analysis even in data-scarce scenarios.
-Contributions. We summarize following contributions:
-    • LLM-as-CTI-Copilot: We propose CyLens, a scalable and adaptive system for CTI practice.
-CyLens is designed to accommodate diverse needs, making it well-suited for both research and industry applications.
-    • Empirical Insights: Through extensive evaluations, we not only demonstrate CyLens’s effectiveness, but also reveal the varying performance of SOTA and industry-leading LLMs, highlighting the inherent complexity of CTI analysis and difficulty of achieving consistent expertise.
-    • Blueprint Perspective: The design and development of CyLens serve as a blueprint for integrating LLM copilots into cybersecurity practice.
-This endeavor not only advances the capabilities of CTI systems but also provides insights that could drive the development of expert-level AI across other high-stakes domains [^1] Codes and datasets are organized at https://anonymous.4open.science/r/security-agent-4BB7/.
-2 Background and Related Work
-2.1 Cyber Threat Ingelligence (CTI)
-AI-for-Security has been a long-standing research area, with extensive studies covering a broad range of cybersecurity topics, including anomaly detection [109, 68], malware classification[47, 69], and intrusion detection[48, 31].
-Among these, CTI has emerged as a systematic approach to threat management, enabling organizations to collect, analyze, and operationalize intelligence on cyber threats [20, 95, 89].
-Although there is no universal specification for CTI practices, we categorize CTI into six critical threat management tasks based on extensive literature [82, 32, 86, 54, 42, 75, 89, 71, 114, 43] and the MITRE initiative [3], a foundational effort in the field.
-As illustrated in Figure 1-(a), the lifecycle of threat intelligence begins with the extraction and analysis of threat properties (❶ & ❷) [42, 105, 75], progresses to the identification of adversary community (❸ & ❹) [32, 86, 54, 114], and culminates in the development of effective countermeasures (❺ & ❻) [89, 71, 43].
-In this work, we set the scope of the CTI tasks as follows:
-    • Attribution–Identifying the origin of cyber threats by profiling threat actors, analyzing tactics, techniques, and procedures (TTPs), and tracing exploitation paths to link cyber incidents to specific adversaries or groups.
-Attribution helps organizations understand attacker intent, capabilities, and historical attack patterns [105, 42].
-    • Contextualization–Enriching raw threat intelligence with relevant metadata, such as affected systems, attack infrastructures, and potential impacts.
-Contextualization enables security analysts to assess threats with a comprehensive picture of the threat landscape[52, 75].
-    • Detection–Identifying cyber threats using available indicators, such as CVE [18] and CWE [16]. Detection assists the recognition of both known and emerging threats [114, 32].
-    • Correlation–Connecting disparate threats to uncover hidden relationships between attack campaigns and disclose the broader landscape of adversarial community.
-Correlation enhances situational awareness and strengthens incident investigations by providing a more comprehensive view of threat activities [86, 54, 52].
-    • Prioritization– Ranking threats based on severity, exploitability, and potential impact to determine the urgency of remediation.
-Effective prioritization optimizes resource allocation by focusing on the most critical threats [89, 43].
+Lifecycle-Oriented Threat Analysis – CTI practice encompasses a lifecycle of analyzing and counteracting cyber threats (Irshad and Siddiqui, 2023; Xiang et al., 2024; Qiang et al., 2017; Sun et al., 2023; Patil and Malla, 2022).
+As illustrated in Figure 2, CyLens is aligned with this lifecycle and supports five groups of analytical tasks.
+To investigate the nature and origin of threats, CyLens enables ❶ attribution by identifying threat actors and collecting technical evidence, such as TTPs and campaign, to trace the source of an attack.
+It further facilitates ❷ contextualization by enriching metadata, including affected systems and infrastructures, to provide a comprehensive understanding of the threat landscape.
+Next, ❸ correlation involves linking known identifiers, such as CVEs (Corporation, nd), while also uncovering adversarial networks by identifying related or concurrent threats.
+CyLens supports ❹ prioritization by assessing the severity of threats and estimating the likelihood of exploitation, enabling organizations to allocate resources efficiently.
+Finally, based on these analytical foundations, ❺ remediation supports the development of actionable countermeasures to effectively neutralize threats and reduce their broader impact.
+Scalable Knowledge InfusionCyLens incorporates extensive CTI knowledge from a diverse set of threat databases and repositories, including MITRE-CVE (Corporation, nd), CWE (Corporation, 2006), CAPEC (Corporation, 2007), ATT&CK (Corporation, 2013), NVD (National Institute of Standards and Technology, 2025), Exploit-DB (Offensive Security, 2025), and their correlated threat reporting platforms (e.g., IBM-XForce (IBM X-Force, 2025), Apache (Apache Software Foundation, 2025a), and Red Hat (Red Hat, Inc., 2025)), resulting in a collection of 271,570 threat reports.
+To integrate extensive CTI knowledge, CyLens adopts a curriculum training methodology, where the “curriculum” consists of ordered threat reports designed to progressively embed CTI expertise into the model.
+This approach enables scalable knowledge infusion by allowing the training process to incorporate both extensive historical data and emerging threats.
+By adjusting pacing steps and balancing newer data with earlier content, curriculum training eliminates the need for retraining from scratch.
+Adapting through TuningCyLens exhibits strong adaptability in addressing specific analytical needs, such as threat analysis related to particular products, vendors, or threat groups.
+To enable this, we apply instructional fine-tuning (or simply instruction tuning), which leverages task-oriented datasets aligned with critical CTI tasks (❶–❺).
+Within these datasets, we introduce a cascading reasoning framework that reflects the inherent task dependencies across the CTI lifecycle.
+This structured design allows CyLens to emulate best practices as expert-level CTI analysis.
+Modularized Inference – At inference time, we integrate specialized design modules to enhance CyLens’s capacity for rigorous and reliable reasoning.
+For example, threat actor attribution involves applying named entity recognition (NER) to threat reports and retrieving relevant, up-to-date metadata regarding recent exploitation activities.
+By systematically organizing CyLens ’s inference process through these modules, we improve the consistency and precision of its analytical outputs.
+Validated Proficiency – In experiments, we collect real-world threat reports from platforms including Oracle Security Alerts (Oracle Corporation, 2025), Red Hat Bugzilla (Red Hat, Inc., 2025), Adobe Security Bulletin (Adobe Systems Incorporated, 2025), and the Apache Mailing List (Apache Software Foundation, 2025a).
+We design two groups of experiments to evaluate the scalability and adaptability of CyLens.
+For scalability evaluation, we evaluate CyLens’s effectiveness across a large volume of threats spanning CTI tasks ❶–❺, considering multiple analytical targets under each task.
+For example, threat contextualization includes identifying affected systems, attack infrastructure, and potential impacts.
+For adaptability evaluation, we focus on specialized threats targeting specific products (e.g., Acrobat), vendors (e.g., Adobe), or threat categories (e.g., DDoS).
+These scenarios reflect the diverse and evolving needs of industrial organizations and research institutions (Saeed et al., 2023; Van Haastrecht et al., 2021; Sakellariou et al., 2022).
+In both evaluations, we categorize threats into two groups: (i) Historical Threats—previously recorded threats that security analysts may not be fully aware of; and (ii) Zero-Day Threats—newly emerging threats that have not yet been documented.
+Across both categories, CyLens consistently outperforms state-of-the-art (SOTA) models and industry-leading LLMs (e.g., ChatGPT and Gemini), demonstrating its versatility in addressing a broad range of cybersecurity challenges and its effectiveness in supporting proactive threat hunting within a dynamic threat landscape.
+Contributions. We summarize the following contributions:
+    • We propose CyLens, a scalable and adaptive agentic system for CTI.
+CyLens supports lifecycle-oriented threat analysis, encompassing tasks from attribution to remediation.
+To accommodate varying computational constraints, CyLens is implemented in three model sizes: 1B, 8B, and 70B.
+Through comprehensive knowledge infusion, followed by instruction tuning and modularized inference, CyLens is equipped with deep CTI expertise and consistently delivers rigorous, reasoning-driven analysis.
+    • Through extensive evaluations on historical and zero-day threats, we not only demonstrate CyLens’s effectiveness, but also reveal the varying performance of SOTA and industry-leading LLMs, highlighting the inherent complexity of CTI analysis and difficulty of achieving consistent expertise.
+    • The design and development of CyLens serve as a blueprint for integrating agentic LLM into cybersecurity practice.
+This endeavor not only advances the capabilities of CTI systems but also provides insights that could drive the development of expert-level AI across other high-stakes domains.
+[^1] Codes and datasets are organized at https://anonymous.4open.science/r/security-agent-4BB7/
+2. Background and Related Work
+2.1. Cyber Threat Ingelligence (CTI)
+AI-for-Security has been a long-standing research area, with extensive studies covering a broad range of cybersecurity topics, including anomaly detection (Yaseen, 2023; Pang et al., 2021), malware classification(Kalash et al., 2018; Pascanu et al., 2015), and intrusion detection(Khraisat et al., 2019; Frank, 1994).
+Among these, CTI has emerged as a systematic approach in cybersecurity, enabling organizations to collect, analyze, and operationalize cyber threats (Cybersecurity and , CISA; Wagner et al., 2019; Sun et al., 2023).
+Although there is no universal taxonomy in CTI, we categorize its involved tasks into five critical aspects based on extensive literature (Gao et al., 2020; Lin et al., 2023; Irshad and Siddiqui, 2023; Qiang et al., 2017; Sun et al., 2023; Patil and Malla, 2022; Zhu et al., 2024b; Jacobs et al., 2020; Kouremetis et al., 2025) and the MITRE initiative (Al-Sada et al., 2024), a foundational effort in the field.
+As illustrated in Figure 2, the lifecycle of threat intelligence begins with investigatig threat properties (❶ & ❷) (Irshad and Siddiqui, 2023; Xiang et al., 2024; Qiang et al., 2017), progresses to identify adversary community (❸) (Gao et al., 2020; Lin et al., 2023; Zhu et al., 2024b), and culminates in developing countermeasures (❹ & ❺) (Sun et al., 2023; Patil and Malla, 2022; Jacobs et al., 2020).
+Formally, this work set the scope of the CTI tasks as follows:
+    • Attribution–Identifying the origin of cyber threats by profiling threat actors, analyzing tactics, techniques, and procedures (TTPs), and tracing attack campaign to link cyber incidents to specific adversaries or groups.
+Attribution helps organizations understand attacker intent, capabilities, and historical attack patterns (Xiang et al., 2024; Irshad and Siddiqui, 2023).
+    • Contextualization–Enriching relevant metadata, such as affected systems, attack infrastructures, and potential impacts.
+Contextualization enables security analysts to assess threats with a comprehensive picture of the threat landscape(Li et al., 2022; Qiang et al., 2017).
+    • Correlation–Linking available threat identifiers, such as CVE (Corporation, 2013) and CWE (Corporation, 2006), and uncovering related or concurrent activities within the same attack campaigns.
+Correlation supports to identify both known and emerging threats (Zhu et al., 2024b; Gao et al., 2020), while exposing the broader landscape of the attack community (Lin et al., 2023).
+    • Prioritization– Ranking threats based on severity and likelihood of exploitation to determine their remediation urgency.
+Effective prioritization optimizes resource allocation by focusing on the most critical and high-risk threats.(Sun et al., 2023; Jacobs et al., 2020).
     • Remediation–Developing and recommending actionable countermeasures, including security patches, mitigation strategies, and responsible advisories.
-Remediation ensures that organizations can effectively neutralize threats and reinforce their security posture[43, 74].
+Remediation ensures that organizations can effectively neutralize threats and reinforce their security posture (Jacobs et al., 2020; Poehlmann et al., 2021).
 Within each CTI task, we define distinct analytical targets.
-For instance, prioritization involves assessing both severity, measured by the Common Vulnerability Scoring System (CVSS) [60], and exploitation potential, measured by the Exploit Prediction Scoring System (EPSS) [29].
+For instance, prioritization involves assessing both severity, measured by the Common Vulnerability Scoring System (CVSS) (National Institute of Standards and Technology (2025), NIST), and exploitation potential, measured by the Exploit Prediction Scoring System (EPSS) (FIRST.org, 2025).
 The detailed analytical targets for each CTI task are presented in Table 2 and are evaluated extensively in section 5.
-2.2 LLM-as-Copilot
-Figure 2: Development pipeline of LLM copilots.
-Due to their vast inherent knowledge and ability to handle large-scale downstream tasks, LLMs have been extensively adapted for domain-specific applications, an effort referred to as “LLM-as-Copilot” [51, 34, 90].
-Unlike the broader concept of an "agent," which is designed to perform human-like behaviors, a copilot specifically emphasizes its role as a human assistant [85, 55, 88].
-This distinction is particularly critical in high-stakes domains such as cybersecurity [112, 28], healthcare [4, 98, 108], and finance [107, 111, 70], where expert guidance and decision support are essential.
-Training Methodology.
-As illustrated in Figure 2, the development of LLM copilots follows a pre-training and fine-tuning trajectory.
-Pre-training leverages large-scale corpora to embed extensive domain expertise into base LLMs through a causal language modeling objective [9, 25, 45, 27].
-Subsequently, fine-tuning incorporates environment-aware instructions and utilizes supervised learning [61, 78] to refine the model’s reasoning capabilities for practical applications.
-This pre-training followed by fine-tuning approach has proven particularly effective in agent development requiring broad domain expertise, such as existing efforts in AI-assisted programming [2, 91].
-Similarly, the interconnected nature of cyber threats necessitates a deep and expansive CTI knowledge.
-Thus, this work adopts the same development pipeline to enhance LLM capabilities, advancing them toward the creation of an effective and practical cybersecurity copilot.
-Prompting.
-Besides training methodologies, inference-time prompting plays a crucial role in regulating an LLM’s behavior during execution [63, 36].
-While developing LLM copilots cannot rely solely on prompting, it provides flexibility by enabling them to interact with real-world resources, such as functional tools [87] or retrieving web-based knowledge [53].
-In this work, we enhance CyLens’s inference by implementing 6 NLP modules, integrated through either additional functions or strategic prompting techniques.
-These enhancements improve CyLens’s reasoning, contextual awareness, and overall effectiveness in CTI analysis.
-3 Data Management and Pre-Training
-In this section, we describe the collection and preprocessing of threat reports used in pre-training (section 3.1) and introduce a curriculum-based pre-training approach that integrates large-scale cyber threat knowledge into CyLens (section 3.2).
-The subsequent section (section 4) details the fine-tuning and inference design.
-3.1 Threat Data: Collection and Preprocessing
+2.2. Agentic LLM
+As extensively studied in recent works, the agentic LLM paradigm follows a two-stage development process: $\theta_{\texttt{pt}}\rightarrow\theta_{\texttt{agent}}$ (Xu et al., 2023; Xi et al., 2023; Yang et al., 2024), where a pre-trained LLM $\theta_{\texttt{pt}}$ serves as the foundation due to its inherent language comprehension and reasoning abilities.
+The model is then customized into $\theta_{\texttt{agent}}$ to enhance its domain-specific expertise.
+This customization may involve re-training to infuse large-scale domain knowledge into LLMs—such as in programming synthesis tasks (Xu et al., 2023)—or instruction tuning to adapt LLMs to highly specialized applications, such as cardiological diagnostics (Zhou et al., 2024).
+During inference, agentic LLMs can be further enhanced using specialized prompts to structure to regularize their reasoning processes (Wei et al., 2022; Hu et al., 2024), or by integrating tool functions (e.g., APIs) to augment their problem-solving capabilities in domain-specific tasks (Qin et al., 2023a, b).
+For example, given a threat report $d$, one can implement a named entity recognition function $\texttt{NER}(\cdot)$ callable by the LLM, such that $\texttt{NER}(d)\rightarrow\mathcal{E}$ extracts threat-related entities (evidence) $\mathcal{E}$ from $d$.
+These tool functions may be implemented as embodied interfaces triggered through prompt instructions (Schick et al., 2023; Li et al., 2024).
+3. Methodology
+3.1. Overview
+This section details the design of CyLens. Following the agentic LLM paradigm (section 2.2), CyLens requires both customization and tailored inference to address heterogeneous CTI tasks.
+Given the diverse analytical requirements and task-specific disparities in CTI, we adopt a two-stage customization consisting of re-training and instruction tuning to enhance LLMs with CTI expertise.
+In the re-training stage, we build a large-scale threat corpus (section 3.2) and develop a curriculum training approach (section 3.3) to progressively infuse CTI knowledge into LLMs.
+Next, we apply instruction tuning using a dataset curated with a cascading reasoning layout (section 3.4), which reflects the dependencies among CTI tasks.
+A fine-tuned LLM can then be guided to follow this reasoning workflow, enabling it to perform multi-step threat analyses in a structured and expert-like manner.
+During inference, CyLens uses six NLP modules to structure its generation workflow (section 3.5).
+These include topic modeling to determine the query intent, entity and relation extraction, information retrieval, and finally, reasoning and evidence summarization to generate coherent, task-specific responses.
+Together, these modules help CyLens produce accurate, structured outputs while minimizing hallucinations during open-minded generation.
+3.2. Threat Corups Construction
 Raw Data Collection.
-Given that threat data is predominantly represented by CVEs [100, 102], we focus on collecting CVE-centered metadata from several vulnerability databases: MITRE-CVE [14], NVD [59], Exploit-DB [62], CWE [16], CAPEC [17], and ATT&CK [18], including (but not limited to) associated threat actors, TTPs, affected systems, attack infrastructures, and strategical mitigation approaches.
-To enrich the collected cyber threat knowledge, we further extract reports and evidence from third-party references such as GitHub commits, Apache [5], and IBM X-Force [39], incorporating diverse cybersecurity activities and countermeasure strategies.
-These additional sources enhance the comprehensiveness of our dataset by providing real-world exploit details and remediation insights.
+Given that threat data is predominantly represented by CVEs (Wikipedia contributors, 2025a, c), we focus on collecting CVE-centered metadata from several threat databases: MITRE-CVE (Corporation, nd), NVD (National Institute of Standards and Technology, 2025), Exploit-DB (Offensive Security, 2025), CWE (Corporation, 2006), CAPEC (Corporation, 2007), and ATT&CK (Corporation, 2013).
+To enrich the collected cyber threat knowledge, we further extract reports and evidence from third-party references such as Oracle, Red Hat Bugzilla, Adobe Security Bulletin, and the Apache Mailing List, incorporating diverse cybersecurity activities and countermeasure strategies.
 Table 10 outlines the data sources and the corresponding metadata collected.
-Pre-Training Corpus.
-Next, we preprocess the collected metadata to construct CVE-centric threat reports, which are used as pre-training corpus.
-As outlined in Algorithm 1, we employ a combination of three approaches to ensure the generation of high-quality reports and diverse layouts:
-    • In-Context Learning: We extract sample layouts and phrasing styles from threat reports published by Microsoft [13], Cisco [41], CISA [20], Europol [26], and ENISA [30].
-These examples are used as demonstrations $\mathcal{S}$ to guide the generation.
+Corpus Generation.
+Next, we curate the collected metadata to construct CVE-centric threat reports, which are then aggregated as threat corpus used for re-training.
+As outlined in Algorithm 1, we employ a combination of three approaches to ensure the generation of high-quality threat reports and diverse layouts:
+    • In-Context Learning: We extract sample layouts and phrasing styles from threat reports published by Microsoft (Corporation, 2024), Cisco (Intelligence, 2024), CISA (Cybersecurity and , CISA), Europol (Europol, 2017), and ENISA (for Cybersecurity , ENISA).
+These examples (or “demos”) $\mathcal{S}$ are used to guide the generation.
 For each generation, we select an example report $s\in\mathcal{S}$ and construct a 1-shot prompt, designed to replicate the specific style and layout of $s$.
@@
 Subsequently, a different LLM is randomly chosen from $\Theta$ to revise the generated reports, introducing variability and enhancing quality by reducing single-model bias.
-    • Output Sampling: We adjust two decoding hyperparameters—temperature and top-p—which introduce controlled variability in the outputs.
-This approach, as studied in [11, 72, 115], helps balance randomness and coherence, ensuring diverse yet meaningful generations.
+    • Output Sampling: We adjust two decoding hyperparameters, temperature and top-p, which introduce controlled variability in the outputs, thus helps balance randomness and coherence, ensuring diverse yet meaningful generations (Chen et al., 2021; Pearce et al., 2023; Zhu et al., 2024a).
 Input: $\Theta$ – a set of LLMs; $\mathcal{M}$ – collected metadata; $\mathcal{S}$ – example threat reports; $\mathcal{P}_{\texttt{gen}}$ – generation prompt; $\mathcal{P}_{\texttt{rev}}$ – revision prompt;
@@
 3       1-shot demo: $s\leftarrow\texttt{RandSelectFrom}(S)$ ;
-4       Input prompt: $\mathcal{P}_{m}\leftarrow\mathcal{P}\oplus s\oplus m$;
+4       Input prompt: $\mathcal{P}_{m}\leftarrow\mathcal{P}_{\texttt{gen}}\oplus s\oplus m$;
        // Report Generation and Revision
@@
 Algorithm 1 Corpus Generation
+Table 1. Base LLMs and re-training settings. “Cp” – Corpus; “Tk” – token. We assign different data volumes each model given their capacities and optimization requirements.
+CyLens Base LLM Curriculum Training
+# Cp # Tk $T$ $T_{1}$ $T_{2}$ $\beta$
+CyLens-1B Llama-3.2-1B 28K 0.27B 4 3 3 1.0
+CyLens-8B Llama-3.1-8B 115K 1.1B 10 5 5 1.0
+CyLens-70B Llama-3.3-70B 271K 2.5B 20 10 10 1.0
 Due to space constraints, we defer the example reports, input prompts, and the corpus scale to Appendix A.2.
-Table 1: Details of the fine-tuning dataset.
+Table 2.
+Details of CTI tasks.
 Each task targets a specific aspect of CTI analysis, incorporating varying dependencies from prior tasks, and involving different NLP modules: TOM–Topic Modeling; NER–Named Entity Recognition; REL–Relation Extraction; RAG–Retrieval-Augmented Generation (using web resources); REA–(any type of) Reasoning; SUM–Summarization.
+NLP modules are integrated at inference time to guide CyLens ’s behavior.
 CTI Task Analytical Target Dependency NLP Module
 TOM NER REL RAG REA SUM
-❶ Attribution Threat Actor, TTPs, Exploitation Path Initial Evidence
+❶ Attribution Threat Actor, TTPs, Campaign Initial Evidence
 ❷ Contextualization Affected System, Attack Infra, Impact Initial Evidence
-❸ Detection CVE, CWE Identifier Initial Evidence+❶❷
-❹ Correlation Related CVE Identifier Initial Evidence+❶❷❸
-❺ Prioritization Severity (CVSS), Exploitation (EPSS) Initial Evidence+❶❷❸
-❻ Remediation Patch Tool, Code, Mitigation Advisory Initial Evidence+❶❷❸❹❺
-3.2 Curriculum Pre-Training
-Next, we pre-train the base LLMs (detailed in Table 9) to embed large-scale cybersecurity knowledge.
-We design a “curriculum” strategy that progressively introduces the corpus throughout the training process.
+❸ Correlation CVE, CWE Identifier Initial Evidence+❶❷
+❹ Prioritization Severity (CVSS), Exploitation (EPSS) Initial Evidence+❶❷❸
+❺ Remediation Patch Tool, Code, Mitigation, Advisory Initial Evidence+❶❷❸❹
+3.3. Curriculum Training (Knowledge Infusion)
+Next, we re-train base LLMs (detailed in Table 1) to infuse large-scale CTI knowledge. We design a “curriculum” strategy that progressively introduces the corpus throughout the training process.
 Curriculum Design.
@@
 Formally:
-$\textsc{Pace}(D,t)=\begin{cases}D[:|D|\cdot\frac{t}{T_{1}}]&t\leq T_{1}\\ D&T_{1}<t\leq T_{2}\\ D\cup D[\beta\cdot|D|\cdot\frac{t-T_{2}}{T-T_{2}}:]&t>T_{2}\end{cases}$ (1)
+(1) $\textsc{Pace}(D,t)=\begin{cases}D[:|D|\cdot\frac{t}{T_{1}}]&t\leq T_{1}\\ D&T_{1}<t\leq T_{2}\\ D\cup D[\beta\cdot|D|\cdot\frac{t-T_{2}}{T-T_{2}}:]&t>T_{2}\end{cases}$
 The pacing function in Eq.
@@
 This augmentation is balanced by a factor $\beta$.
-Table 9 shows values of $T$, $T_{1}$, and $T_{2}$.
-Pre-Training Objective.
-Identical to the pre-training objective of Llama [92], we employ Causal Language Modeling (CLM), an autoregressive training objective where the LLM predicts the next token in a sequence based on preceding tokens.
+Table 1 shows values of $T$, $T_{1}$, $T_{2}$, and $\beta$.
+Training Objective.
+Identical to the re-training objective of Llama (Touvron et al., 2023), we employ Causal Language Modeling (CLM), an autoregressive training objective where the LLM predicts the next token in a sequence based on preceding tokens.
 Let $x_{j}$ denote a token in a threat report $d$ with $L_{d}$ tokens, we have $d=\{x_{j}\}_{j=1}^{L_{d}}$.
 The CLM objective is formulated as:
-$\ell_{\texttt{CLM}}(\theta,D)=-\mathbb{E}_{d\sim\mathcal{D}}\left[\sum_{j=1}^{L_{d}}\log p_{\theta}(x_{j}|x_{<j})\right]$ (2)
-Here, the model $\theta$ learns to maximize the likelihood of the next token $x_{j}$ given the preceding context $x_{<j}$, across all tokens in a threat report $d$.
-This objective enables the model to effectively internalize and represent the knowledge included in the cyber threat corpus $D$.
-4 Fine-Tuning, Inference, and Customization
-This section details the development of CyLens after pre-training, focusing on task-oriented dataset construction for fine-tuning (section 4.1), modularized inference (section 4.2), and organizational customization (section 4.3).
-4.1 Fine-Tuning with Cascading Reasoning
+(2) $\ell_{\texttt{CLM}}(\theta,D)=-\mathbb{E}_{d\sim\mathcal{D}}\left[\sum_{j=1}^{L_{d}}\log p_{\theta}(x_{j}|x_{<j})\right]+\lambda\|\theta-\theta_{0}\|_{2}^{2}$
+Here, the model $\theta$ is trained to maximize the likelihood of each token $x_{j}$ given its preceding context $x_{<j}$, across all tokens in a threat report $d$.
+To prevent overfitting to the training corpus, we introduce a regularization term $\|\theta-\theta_{0}\|_{2}^{2}$, where $\theta_{0}$ denotes the original pre-trained parameters.
+This regularization encourages the model to retain its inherent language understanding and reasoning capabilities.
+The resulting objective function strikes a balance between effective CTI knowledge infusion and preserving the general-purpose strengths of the base LLM.
+3.4. Instruction Tuning via Cascading Reasoning
 Data Construction.
-CTI analysis follows a cascading process, which starts with an initial set of evidence and progressively gathers additional information to refine the analysis [64, 81].
-To equip LLMs with cascading analytical capabilities, we design the fine-tuning dataset to reflect progressive CTI activities: $D=\{E_{0}\cup(Q_{i},A_{i})_{i=1}^{n}\}$ where $E_{0}$ represents the initial evidence, and each query-answer pair $(Q_{i},A_{i})$ corresponds to a step in the cascading reasoning process:
+CTI analysis follows a cascading process, which starts with an initial set of evidence and progressively gathers additional information to refine the analysis (Oosthoek and Doerr, 2021; Saeed et al., 2023).
+To equip LLMs with cascading analytical capabilities, we design a dataset to reflect progressive CTI activities: $D=\{E_{0}\cup(Q_{i},A_{i})_{i=1}^{n}\}$ where $E_{0}$ represents the initial evidence, and each query-answer pair $(Q_{i},A_{i})$ corresponds to a step in the cascading reasoning process:
 $(A_{1}|E_{0},Q_{1})\rightarrow(A_{2}|E_{0},A_{1},Q_{2})\rightarrow...\rightarrow(A_{n}|E_{0},A_{1},...,A_{n-1},Q_{n})$
-Example 1
+Example 0.
 Below is a cascading reasoning example.
+$\mathbf{E_{0}}$: A security monitoring system detected suspicious outbound traffic from several workstations within the organization’s network.
+Traffic analysis indicates command-and-control (C2) communication with IP addresses known to be associated with APT28, with observed indicators of $\cdots$
+Step 1: Threat Actor Attribution
+$\mathbf{Q_{1}}$: From above, what threat actor is likely responsible?
 […]
-Figure 3: An illustration of modularization in CyLens, using the same threat report as described in Example 1.
-Table 1 presents the analytical targets for CTI tasks ❶-❻, where each task builds upon the preceding ones.
-As illustrated in Example 1, identifying CVEs requires first pinpointing threat actors, which subsequently guides the prioritization.
-To construct the fine-tuning dataset, we aggregate references from MITRE-CVE, NVD, and ATT&CK.
+Step 2: Vulnerability Identification
+$\mathbf{Q_{2}}$: What known vulnerabilities (CVEs) are commonly associated with APT28 in recent campaigns?
+$\mathbf{A_{2}}$: APT28 has exploited CVE-2023-12345 (Windows Kernel Privilege Escalation) and CVE-2023-67890 (Remote Code Execution via malicious DLL sideloading).
+Step 3: Prioritization of Threat Response $\cdots$
+Figure 3. An illustration of modularization in CyLens, using the same threat report as described in Example 3.1.
+[…]
+Modularized inference.
+Table 2 presents the analytical targets for CTI tasks ❶-❺, where each task builds upon the preceding ones.
+As illustrated in Example 3.1, identifying CVEs requires first pinpointing threat actors, which subsequently guides the prioritization.
+To construct the dataset, we aggregate references from threat databases (e.g., MITRE and NVD).
 For each reference, we extract available evidence as listed in Table 2 and utilize descriptive reports as the initial evidence set $E_{0}$.
-We then generate $(Q_{i},A_{i})$ pairs following the format exemplified in Example 1.
-It is important to note that not all constructed data samples encompass the full pipeline of tasks ❶-❻.
-Detailed dataset statistics are provided in Appendix B.
-Fine-Tuning.
+We then generate $(Q_{i},A_{i})$ pairs following the format exemplified in Example 3.1.
+It is important to note that not all constructed data samples encompass the full pipeline of tasks ❶-❺.
+Dataset statistics are provided in Appendix B.
+Instruction Tuning.
 To enable CyLens with cascading reasoning capabilities, we fine-tune it using the dataset $D$ described above.
-The fine-tuning process follows a standard instruction-tuning objective [106, 44], where the model is trained to generate a sequence of answers conditioned on the preceding context.
+This process follows a standard instruction-tuning objective (Xu et al., 2023; Ji et al., 2024), where the model is trained to generate a sequence of answers conditioned on the preceding context.
 Let $\theta$ represent the parameters of CyLens.
 The objective function is defined as:
-$\begin{split}\mathcal{L}(\theta)=-\mathbb{E}_{(E_{0},Q_{1\sim n},A_{1\sim n})\in D}\ell(E_{0},Q_{1\sim n},A_{1\sim n},\theta)\\ \text{s.t.}\quad\ell=\sum_{i=1}^{n}\log P_{\theta}(A_{i}|E_{0},A_{1},\dots,A_{i-1},Q_{i})\end{split}$ (3)
+(3) $\begin{split}\mathcal{L}(\theta)=-\mathbb{E}_{(E_{0},Q_{1\sim n},A_{1\sim n})\in D}\ell(E_{0},Q_{1\sim n},A_{1\sim n},\theta)\\ \text{s.t.}\quad\ell=\sum_{i=1}^{n}\log P_{\theta}(A_{i}|E_{0},A_{1},\dots,A_{i-1},Q_{i})\end{split}$
 where $P_{\theta}(A_{i}|E_{0},A_{1},\dots,A_{i-1},Q_{i})$ represents the probability assigned by the model to the correct response $A_{i}$ given the accumulated context up to step $i$.
-Table 2: Evaluation results for ❶ Attribution tasks, where "Hist."–historical threats, "0-Day"–zero-day threats.
+3.5. Modularized Inference
+CyLens’s inference also follows the cascading reasoning layout.
+Each reasoning step $(E_{0},A_{1},\dots,A_{i-1},Q_{i})\rightarrow A_{i}$ employs various NLP functions to modularize the analytical process.
+Table 2 summarizes the NLP techniques applied to each CTI task.
+These techniques are implemented through the following approaches:
+Topic Modeling.
+We implement topic modeling to interpret users’ intentions regarding the tasks they aim to perform.
+Specifically, we use BERTopic, a transformer-based package that eliminates the need for manually setting hyperparameters, as required in conventional methods (Blei et al., 2003; Lee and Seung, 2000).
+Given a threat report $E_{0}$, we implement a two-step approach: first, based on CTI tasks ❶-❺, and then, based on specific analytical targets, as outlined in Table 2.
+Example 0.
+Topic modeling (Figure 3-(b)).
+Using the same example as in 3.1, the topic modeling step first determines the task as threat attribution based on $Q_{1}$.
+It then further specifies the analytical target as the threat actor within the threat attribution task.
+Named Entity Recognition (NER) and Relation Extraction.
+Next, we extract entities relevant to the identified topics while simultaneously identifying relationships among them.
+We employ the multi-agent generation as described in section 3.2, using few-shot prompting to instruct entity and relation extraction, where domain-specific entities (e.g., APT group) and their relationships are included as demonstrations within the prompt.
+We defer detailed implementation in section C.2 and C.3.
+Example 0.
+NER and relation extraction (Figure 3-(c)).
+Using Example 3.1, the NER identifies key entities ”APT28” and ”C2 communication” within the threat report.
+Simultaneously, relation extraction determines the ”is associated with” relationship between these entities.
+Retrieval Augmented Generation (RAG).
+We implement RAG to retrieve relevant web documents through web screening to enrich threat reports.
+This is especially useful for zero-day threat hunting, where web-scale CTI intelligence—such as reported threat groups, affected products, and potential patches—is continuously updated.
+Then, we prompt GPT-o1 to rank the retrieved documents based on their relevance to the threat report $E_{0}$.
+To control latency, we implement a cache to store retrieved documents.
+Caching documents are indexed by topic modeling results (i.e., CTI tasks and analytical targets) and specific entities (e.g., APT28 or C2 communication).
+Reasoning and Summarization.
+We implement reasoning and summarization through prompting in CyLens, where reasoning is applied when deeper analysis or multi-step thinking is required.
+For example, reasoning is used to analyze severity level based on gathered evidence.
+Summarization is then applied to synthesize insights, tailoring the output to a specific CTI task by leveraging results from topic modeling, NER, relation extraction, and RAG.
+Additional implementation details (e.g., prompts) of the six NLP modules are provided in appendix C.
+Cascading Reasoning in Inference.
+Since CTI tasks exhibit dependencies (Table 2), we minimize latency by concurrently executing topic modeling, NER, relation extraction, and RAG across all prerequisite tasks.
+Reasoning and summarization then proceed sequentially, following task dependencies to generate structured, step-by-step findings and conclusions.
+4. Experimental Setting
+Table 3.
+Evaluation results for ❶ Attribution tasks, where ”Hist.”–historical threats, ”0-Day”–zero-day threats.
 All metrics are scaled to 0-100 (%) for consistency.
 The boldface and underline highlight the best and second-best results, respectively.
-Threat Actor TTP Exploitation Path
+Threat Actor TTP Campaign
 Model Accuracy BERT Score Precision Recall F1 Score IoU Accuracy BERT Score
@@
 CyLens-70B 96.53 88.29 87.93 84.52 89.63 87.57 88.26 85.46 88.94 86.50 80.08 76.21 91.53 86.06 87.32 87.03
-4.2 Modularized Inference
-At each reasoning step $(E_{0},A_{1},\dots,A_{i-1},Q_{i})\rightarrow A_{i}$ during inference, we employ various NLP techniques to modularize the analytical process.
-Table 1 summarizes the NLP techniques used for each task.
-These techniques are implemented through the following approaches:
-Topic Modeling.
-We implement topic modeling to interpret users’ intentions regarding the tasks they aim to perform.
-Specifically, we use BERTopic, a transformer-based package that eliminates the need for manually setting hyperparameters, as required in conventional methods [7, 50].
-Given a threat report $E_{0}$, we implement a two-step approach: first, based on CTI tasks ❶-❻, and then, based on specific analytical targets, as outlined in Table 1.
-Example 2
-Topic modeling (Figure 3-(b)).
-Using the same example as in 1, the topic modeling step first determines the task as threat attribution based on $Q_{1}$.
-It then further specifies the analytical target as the threat actor within the threat attribution task.
-Named Entity Recognition (NER) and Relation Extraction.
-Next, we extract entities relevant to the identified topics while simultaneously identifying relationships among them.
-We employ the multi-agent generation approach described in section 3.1, utilizing few-shot prompting, where task-specific entities and their relationships are included as demonstrations within the prompt.
-This approach facilitates a structured understanding of cybersecurity threats, enhancing the accuracy and contextual relevance of extracted threat intelligence.
-Retrieval Augmented Generation (RAG). We implement RAG to retrieve relevant web documents through web screening to enrich threat reports. This is especially useful for zero-
-Example 3
-NER and relation extraction (Figure 3-(c)).
-Using Example 1, the NER identifies key entities "APT28" and "C2 communication" within the threat report.
-Simultaneously, relation extraction determines the "is associated with" relationship between these entities.
-day threat hunting, where web-scale CTI intelligence—such as reported threat groups, affected products, and potential patches—is continuously updated.
-Then, we prompt GPT-o1 to rank the retrieved documents based on their relevance to the threat report $E_{0}$.
-To control latency, we implement a cache to store retrieved documents.
-Caching documents are indexed by topic modeling results (i.e., CTI tasks and analytical targets) and specific entities (e.g., APT28 or C2 communication).
-Reasoning and Summarization.
-We implement reasoning and summarization through prompting in CyLens, where reasoning is applied when deeper analysis or multi-step thinking is required.
-For example, reasoning is used to analyze severity level (e.g., medium) based on gathered evidence.
-Summarization is then applied to synthesize insights, tailoring the output to a specific CTI task by leveraging results from topic modeling, NER, relation extraction, and RAG.
-Cascading Reasoning in Inference.
-Since CTI tasks exhibit dependencies (Table 1), we minimize latency by concurrently executing topic modeling, NER, relation extraction, and RAG across all prerequisite tasks.
-Reasoning and summarization then proceed sequentially, following task dependencies to generate structured, step-by-step findings and conclusions.
-4.3 Organizational Customization
-Although CyLens is developed using massive CTI knowledge, it also enables customization to meet specific organizational interests on particular threat categories, specific products, or assisting an entire organization.
-Such customization is achieved by further fine-tuning LLMs but specifying the scope of data—for example, concentrating on cyber threats related to a single organization.
-This targeted customization enhances the relevance of CTI in meeting organizational needs.
-We demonstrate CyLens’s potential for tailored customization in section 5.6.
-Through experiments, we highlight its adaptability and flexibility in addressing diverse CTI challenges across different organizational settings.
-5 Evaluation Results
-This section systematically evaluates CyLens’s performance on CTI tasks ❶-❻ with the following settings:
+This section outlines our experimental setup for CTI tasks ❶–❺, designed to evaluate CyLens’s scalability and adaptability. The corresponding results are presented in section 5 and 6, respectively.
+SYSTEM You are a cybersecurity threat analyst specializing in cyber threat intelligence (CTI).
+Your role is to analyze threat reports, ¡Specify Analytical Target¿, and produce structured threat intelligence analysis grounded in established frameworks (e.g., MITRE ATT&CK).
+You will follow a structured analytical pipeline with the following steps: Resolve Task Dependencies: Determine the prerequisite CTI tasks and analytical targets based on the input query.
+Launch NLP Modules: ¡Explain Six NLP Modules¿.
+USER ¡Report Title¿ ¡Report Excerpt¿ ¡User’s Question¿ ASSISTANT The user’s intent pertains to ¡Topic Modeling¿, which aligns with the prerequisite identification of ¡CTI Tasks¿ and corresponding ¡Analytical Targets¿.
+Proceeding with a structured inference pipeline, we will now extract and analyze the following components: ¡Key Entities¿, ¡Entity Relations¿, ¡Retrieved Evidence¿, ¡Sources¿, and ¡Synthetic Reasoning¿.
+Final Response: ¡Summarization¿
+Figure 4.
+Inference Prompt structure.
+For baselines and CyLens, we prompt them to perform cascading reasoning with the NLP modules defined in Table 2.
+For tasks with prior dependencies, we instruct LLMs to first complete the prerequisite tasks before proceeding.
+Scenarios. To evaluate the scalability and adaptability of CyLens, we consider two experimental scenarios:
+    • Broad CTI — Demonstrates CyLens’s scalability in analyzing a wide range of cyber threats across CTI tasks ❶–❺, providing a comprehensive view of the threat landscape.
+    • Specialized CTI — Demonstrates CyLens’s adaptability to specific cyber threat groups, such as those targeting particular products or vendors tailored to business needs, or threats within a specific category to support focused interests (e.g., (Mishra et al., 2022; Balasubramanian et al., 2025)).
+In both scenarios, CyLens undergoes the same re-training using the threat corpus described in Table 1.
+However, instruction tuning differs based on the evaluation focus.
+For the broad CTI setting, we use a diverse instruction tuning dataset covering a wide spectrum of threats, as reported in Table 11.
+In the Specialized CTI setting, we extract targeted subsets of threats from the broader dataset and independently fine-tune CyLens to emphasize its functions on specific threat groups (e.g., those associated with a particular product, vendor, or threat category).
 Dataset.
-To ensure practicality, we directly assess threats from NVD [59], MITRE [14], and Exploit-DB [62], categorizing into two groups: (1) Historical – threats reported between 2015 and 2024.
-(2) Zero-day – newly reported threats from the past two months during our evaluation period in 2025.
-For each threat, we compile source references from [14, 59] into structured reports and prompt LLMs for analysis.
-The prompt structure is illustrated in Figure 4.
-For different tasks, we modify the system prompt and questions accordingly.
-To ensure fairness, no test threats were included in training. Statistics and detailed settings are provided in Appendix C.
+To ensure practical evaluation, we directly collect threats reported from real-world platforms, including Oracle Security Alerts (Oracle Corporation, 2025), Red Hat Bugzilla (Red Hat, Inc., 2025), Adobe Security Bulletin (Adobe Systems Incorporated, 2025), NVD (National Institute of Standards and Technology, 2025), VulDB (VulDB, 2025), Patchstack (Patchstack, 2025), and public mailing lists from Apache (Apache Software Foundation, 2025a), Debian (Debian Project, 2025), openSUSE (openSUSE Project, 2025), and Fedora (Fedora Project, 2025).
+To avoid data leakage, all threats correlated with these evaluation reports are excluded from both the re-training and instruction tuning phases.
+We categorize the evaluation threats into two groups: (1) Historical — Threats disclosed between 2015 and 2024.
+(2) Zero-Day — Newly disclosed threats from the two months immediately preceding our evaluation period in 2025.
+Table 4 summarizes the statistics of the evaluation dataset.
+Table 4. Statistics of evaluation threats (not training) for broad and specialized CTI. ”Mc.”–MacOS, ”Cr.”–Chrome, ”Lk.”–Linux Kernel, ”Ad.”–Adobe, ”Dl.”–Dell, ”Of.”–Overflow
+# Threat Broad Product Vendor Category
+Mc. Cr. Lk. IBM Ad. Dl. Of. DoS XSS
+Historical 26.5K 2.5K 1.7K 4.6K 3.3K 1.8K 0.8K 7.5K 9.6K 12.8K
+Zero-Day 3.6K 141 48 176 81 72 33 407 721 2.6K
 Baseline.
-We primarily consider two groups of baselines: (1) State-of-the-art (SOTA) CTI agents, SevenLLM [44] and Lily-Cyber [84], specifically developed for CTI, and (2) Industry-leading general-purpose LLMs, such as ChatGPT and Gemini, which are widely used.
+We primarily consider two groups of baselines: (1) SOTA CTI agents, SevenLLM (Ji et al., 2024) and Lily-Cyber (Segolily Labs, 2025) and (2) Industry-leading general-purpose LLMs: ChatGPT-4o/o1, Gemini-Pro, Llama-405B, DeepSeek-R1, and Claude-Opus.
 Additionally, we investigate the impact of ablation baselines in section 7.1.
-Metrics. We use different evaluation metrics depending on the specific CTI task. The definitions, strategies, and rationale for using these metrics are provided in Appendix C.2.
-Table 3: Evaluation results for ❷ Contextualization tasks, where "Hist."–historical threats, "0-Day"–zero-day threats.
+Metrics. We use different evaluation metrics depending on the specific CTI task. The definitions, strategies, and rationale for using these metrics are provided in Appendix D.2.
+5. Evaluation Results: Broad CTI
+Table 5.
+Evaluation results for ❷ Contextualization tasks, where ”Hist.”–historical threats, ”0-Day”–zero-day threats.
 All metrics are scaled to 0-100 (%) for consistency.
@@
 CyLens-70B 94.54 88.17 96.04 93.75 95.28 90.87 90.99 83.28 96.37 90.16 97.69 93.18 97.03 91.65 94.22 84.58 87.33 85.84
-5.1 Threat Attribution
-We begin by attributing the origin of cyber threats with three targets: threat actor, TTP, and exploitation path, as outlined in Table 1. Our results yield several research findings (RF):
+5.1. Threat Attribution
+We begin by attributing the origin of cyber threats with three targets: threat actor, TTP, and campaign, as outlined in Table 2. Our results yield several research findings (RF):
 RF1: Specialized development is necessary.
 As shown in Table 3, industry-leading LLMs, despite their broad applicability across diverse tasks, exhibit limited performance in CTI analysis.
-For instance, ChatGPT-4o has achieved only
-SYSTEM You are a cybersecurity threat analyst specializing in cyber threat intelligence (CTI).
-Your task is to assess potential attack tactics, techniques, and procedures (TTPs), and provide an intelligence assessment following structured industry frameworks such as $\cdots$ USER <Report Title> "Emerging APT Group Exploiting Zero-Day Vulnerabilities in Cloud Infrastructure." <Report Excerpt> Recent intelligence reports indicate that an Advanced Persistent Threat (APT) group has been exploiting an unpatched zero-day vulnerability in cloud service environments.
-The group is using spear-phishing emails to gain initial access, deploying a custom malware variant named ‘CloudPhantom,’ and leveraging compromised API keys to move laterally $\cdots$ <Question> Identify the relevant TTPs used in this attack.
-Let’s think step-by-step, first, determine CTI tasks among $\cdots$ ASSISTANT <CTI Task><Analytical Target><Key Entities><Entity Relations><Retrieved Evidence and Source>$\cdots$<Final Answer>
-Figure 4: Prompt example for TTP analysis.
-For both baselines and CyLens, we prompt them to perform step-by-step analysis following the NLP modules defined in Table 1.
-For tasks with prior dependencies (e.g., remediation), we instruct LLMs to first complete the prerequisite tasks before proceeding.
-40.31% accuracy in threat actor attribution, significantly lower than CyLens-1B (79.52%) and CyLens-8B (83.74%) for the same task.
-This gap underscores the necessity of specialized development, as detailed in section 3 and section 4.
+For instance, ChatGPT-4o has achieved only 40.31% accuracy in threat actor attribution, significantly lower than CyLens-1B (79.52%) and CyLens-8B (83.74%) for the same task.
+This gap underscores the necessity of specialized development, as detailed in section 3.
 Tailoring models specifically for CTI significantly reduces parameter requirements—CyLens-1B and 8B are far more lightweight compared to the 500B–1T parameters of GPT-4o.
@@
 Interestingly, scaling up model size does not always translate to improved CTI effectiveness.
-For instance, CyLens-8B achieves 93.74% recall in TTP attribution, outperforming CyLens-70B (88.26%), suggesting diminishing returns despite undergoing the same development process.
-This indicates that larger models do not alaways yield better performance in CTI tasks, emphasizing the importance of domain-specific optimizations over sheer model size.
+For instance, CyLens-8B achieves 93.74% recall in TTP attribution, outperforming CyLens-70B (88.26%), suggesting diminishing benefits.
+This indicates the importance of domain-specific optimizations instead of sheer model sizes.
 RF3: Task-specific capability varies across baselines.
 As shown in Table 3, baseline models exhibit varying levels of effectiveness across different CTI tasks.
-For instance, GPT-o1 performs obviously worse in attributing threat actors compared to exploitation paths.
+For instance, GPT-o1 performs obviously worse in attributing threat actors compared to campaign.
 A deeper case study reveals that baseline LLMs struggle to accurately interpret CTI-specific terminology related to threat actors.
 For example:
-Case Study 1
+Case Study 1.
 When analyzing Log4Shell (a known vulnerability) with Mirai (a botnet) as the threat actor, GPT-o1 describes the attacker as an “outside attacker”, failing to precisely identify Mirai.
 This demonstrates a fundamental gap in attribution accuracy for complex cyber threats.
-On the other hand, GPT-o1 achieves $>$78% accuracy in analyzing exploitation paths, as this task primarily involves summarizing existing evidence instead of complex reasoning.
-Case Study 2
-In cases where specificity is less critical, a rough description—such as "JNDI lookup and remote code fetching"—sufficiently summarizes the Log4Shell exploitation path.
-While a more precise description would be: "The Log4j library contacts the attacker’s malicious LDAP server, which then responds with a reference to a remote Java class file containing malicious code," the broader summary still effectively conveys the core mechanism.
+On the other hand, GPT-o1 achieves $>$78% accuracy in analyzing attack campaign, as this task primarily involves summarizing existing evidence instead of complex reasoning.
+Case Study 2.
+In cases where specificity is less critical, a rough description—such as ”JNDI lookup and remote code fetching”—sufficiently summarizes the campaign of Log4Shell.
+While a more precise description would be: ”The Log4j library contacts the attacker’s malicious LDAP server, which then responds with a reference to a remote Java class file containing malicious code,” the broader summary still effectively conveys the core mechanism.
 This indicates that baseline models perform better in tasks with higher generation flexibility but face challenges in tasks requiring precise CTI expertise and structured analysis.
-RF4: CyLens Performs consistently across analyses.
-Unlike baseline models, CyLens maintains consistent performance across different analytical targets, highlighting the importance of incorporating comprehensive expertise during development.
-This balanced capability ensures that CyLens does not exhibit biased performance or knowledge gaps throughout the CTI lifecycle.
-The same conclusion is further reinforced by subsequent evaluations.
-5.2 Threat Contextualization
-We continue evaluating CyLens in contextualizing affected systems, attack infrastructure (e.g., domains and servers), and impact.
-Besides the similar findings as outlined in RF1-RF4, our evaluations derive complementary research findings:
+5.2. Threat Contextualization
+We continue evaluating CyLens in contextualizing affected systems, attack infrastructure (e.g., domains and servers), and impact.Our evaluations derive complementary research findings:
+RF4: Larger LLMs tend to be wordy.
+From Table 5, we observe that larger LLMs tend to achieve lower text similarity scores as measured by BERTScore.
+For example, CyLens-70B performs worse than CyLens-8B in generating threat impact descriptions on historical threats (87.33 vs.
+90.03).
+Through detailed case studies, we find that larger LLMs are more likely to produce verbose responses:
+Case Study 3.
+CyLens-8B summarizes an impact analysis concisely: ”The March 2022 cyberattack by the Lazarus Group resulted in the theft of \$620 million from the Ronin Network, significantly impacting the cryptocurrency market and…” This closely aligns with the reference (Wikipedia contributors, 2025b), maintaining factual accuracy while remaining succinct.
+In contrast, CyLens-70B generates a more verbose response: ”In March 2022, the notorious Lazarus Group orchestrated a sophisticated cyber heist, infiltrating the Ronin Network’s defenses to abscond with a staggering \$620 million in cryptocurrency.
+This audacious act not only sent shockwaves through the digital currency markets, causing a temporary plunge in asset values, but also undermined investor confidence…” The latter output is overly wordy and includes sentiment-heavy language (e.g., ”notorious,” ”sophisticated,” ”staggering,” ”audacious”), making it less aligned with the original reference.
+This finding aligns with RF2, reinforcing that larger models are not always more effective for CTI tasks.
+However, to balance overall performance across all CTI tasks, a systematic evaluation is necessary to determine the optimal model scale.
 RF5: Zero-day threat hunting is not always harder.
-Both Table 2 and 3 reveal that analyzing zero-day threats does not always result in worse performance compared to historical threats.
+Both Table 3 and 5 reveal that analyzing zero-day threats does not always result in worse performance compared to historical threats.
 For instance, when assessing the impact of cyber threats, LLMs such as CyLens-1B and SevenLLM sometimes achieve slightly better relevance in their generated analyses.
@@
 However, for tasks such as identifying affected systems, attack infrastructures, or threat actors, zero-day threats often introduce novel and previously unseen products or malware, making analysis more challenging compared to historical threats.
-RF6: Larger LLMs tend to be wordy. Through detailed case study, we observe that larger LLMs tend to generate verbose responses, leading to lower text similarity as measured by BERT score. For example:
-Case Study 3
-CyLens-8B summarizes an impact analysis concisely: "The March 2022 cyberattack by the Lazarus Group resulted in the theft of \$620 million from the Ronin Network, significantly impacting the cryptocurrency market and…" This closely aligns with the reference [101], maintaining factual accuracy while remaining succinct.
-In contrast, CyLens-70B generates a more verbose response: "In March 2022, the notorious Lazarus Group orchestrated a sophisticated cyber heist, infiltrating the Ronin Network’s defenses to abscond with a staggering \$620 million in cryptocurrency.
-This audacious act not only sent shockwaves through the digital currency markets, causing a temporary plunge in asset values, but also undermined investor confidence…" The latter output is overly wordy and includes sentiment-heavy language (e.g., "notorious," "sophisticated," "staggering," "audacious"), making it less aligned with the original reference.
-This finding aligns with RF2, reinforcing that larger models are not always more effective for CTI tasks.
-However, to balance overall performance across all CTI tasks, a systematic evaluation is necessary to determine the optimal model scale.
-RF7: CTI agents should align with real-world practice.
+RF6: CTI agents should align with real-world practice.
 Despite the strong performance of SOTA agents (SevenLLM and Lily-Cyber) on certain tasks (e.g., analyzing TTPs), they struggle with others, such as attack infrastructure analysis.
 We attribute this to a misalignment between their training focus and CTI applications.
-For instance, SevenLLM is trained with a broad emphasis on NLP tasks (e.g., NER), while Lily-Cyber is developed using hand-crafted cybersecurity and hacking-related data pairs [84].
+For instance, SevenLLM is trained with a broad emphasis on NLP tasks (e.g., NER), while Lily-Cyber is developed using hand-crafted cybersecurity and hacking-related data pairs (Segolily Labs, 2025).
 However, these training paradigms often conflate TTPs and threat actors, rather than treating them as distinct analytical components.
 This lack of granularity can hinder precise CTI reasoning, particularly for tasks that require structured analysis of attack properties.
-Although CTI lacks a universally standardized taxonomy, structuring tasks in alignment with real-world CTI databases (e.g., MITRE-CVE and NVD) provides a more practical and reliable foundation for cyber threat analysis.
-This highlights the necessity of LLMs specifically designed for CTI, rather than relying on models with indirect exposure to cybersecurity knowledge or misaligned customization that does not fully reflect real-world CTI practices.
-Table 4: Evaluating ❺ Prioritization performance using CVSS 3.x metrics [60]: AV–Attack Vector, AC–Attack Complexity, PR–Privileges Required, UI–User Interaction, S–Scope, C–Confidentiality Impact, I–Integrity Impact, A–Availability Impact, Base–CVSS Base Severity.
+Table 6.
+Evaluating ❺ Prioritization performance using CVSS 3.x metrics (National Institute of Standards and Technology (2025), NIST): AV–Attack Vector, AC–Attack Complexity, PR–Privileges Required, UI–User Interaction, S–Scope, C–Confidentiality Impact, I–Integrity Impact, A–Availability Impact, Base–CVSS Base Severity.
 We report the accuracy of correctly categorizing each metric (e.g., Attack Vector: Network).
@@
 C70 94.33 100.0 97.56 97.46 98.15 94.74 92.26 95.83 97.57 92.69 100.0 97.71 95.48 93.26 90.97 94.53 92.66 93.16
-5.3 Detection and Correlation
-We now evaluate CyLens and baselines on ❸ detection and ❹ correlation tasks.
-In detection, the objective is to evaluate how effectively LLMs can identify relevant threat identifiers (CVEs and CWEs) from a given threat report.
-This step is crucial for enabling subsequent CTI activities, such as threat correlation and remediation.
-In correlation, we extend the evaluation by examining how LLMs associate other threats (CVEs) using different strategies: (i) Direct correlation–directly correlate additional CVEs.
-(ii) CVEs as intermediates–first detect CVEs from threat reports (from ❸) and then use them for correlation.
-(iii) CWEs as intermediates–first identifying CWEs (i.e., weakness categories, from ❸) and using them for correlations.
-The latter two strategies leverage intermediate evidence (CVEs or CWEs identified in the detection step) to uncover the threat network.
-Our evaluations yield the following new research findings:
-Figure 5: Evaluation results on ❸ Detection tasks using historical threats. The presentation order of LLM names (top-down) follows the same sequence as in Tables 2 (top-down).
-RF8: CyLens demonstrates strong detection relevance.
-Figure 5-(a)(b) shows the hit ratio for the top 10 detected threat identifiers (CVEs, CWEs).
-Identifying CVEs is more challenging due to their vast number ($>$200k) compared to CWEs ($\sim$900).
-However, we also observe that CyLens  is able to detect more relevant CVEs such as those that belong to the same category or affect the same products.
+5.3. Threat Correlation
+We now evaluate CyLens on ❸ correlation tasks.
+The objective is to measure how effectively CyLens can identify relevant threat identifiers (CVEs and CWEs) from a given threat report.
+Moreover, we extend the correlation by identifying additional threats (CVEs) that are not directly from threat reports but are related or concurrent, i.e., the Expansive Correlation.
+We use different strategies: (i) Directly conduct expansive correlation.
+(ii) First find CVEs from threat reports and then expand to additional CVEs.
+(iii) First find CWEs from threat reports and then expand to additional CVEs.
+Our evaluations yield the following findings:
+Figure 5.
+Evaluation results on ❸ Correlation using historical threats (we do not evaluate on 0-day threats as they have incomplete correlation reports).
+The presentation order of LLM names (top-down) follows the same presentation sequence as in Tables 3 (top-down).
+RF7: CyLens demonstrates strong correlation relevance.
+Figure 5-(a)(b) shows the hit ratio for the top 10 correlated threat identifiers (CVEs, CWEs).
+Correlating CVEs is more challenging due to their vast number ($>$200k) compared to CWEs ($\sim$900).
+However, we also observe that CyLens  is able to correlate more relevant CVEs such as those that belong to the same category or affect the same products.
 For example:
-Case Study 4
-Given the threat report: "…Apache HTTP Server allows remote attackers to execute arbitrary code.
-The issue arises from improper handling of user-supplied input in HTTP request headers…" corresponding to CVE-2021-42013.
-LLMs present different detections: GPT-o1 and DeepSeek-R1 incorrectly output CVE-2020-0609 and CVE-2022-2296, both of which are unrelated to Remote Code Execution or Apache.
+Case Study 4.
+Given the threat report: ”…Apache HTTP Server allows remote attackers to execute arbitrary code.
+The issue arises from improper handling of user-supplied input in HTTP request…” is related to CVE-2021-42013.
+LLMs present different outputs: GPT-o1 and DeepSeek-R1 incorrectly output CVE-2020-0609 and CVE-2022-2296, both of which are unrelated to Remote Code Execution or Apache.
 CyLens-70B gives CVE-2023-25690, a highly relevant HTTP Request Smuggling attack on the Apache Server.
-From the above case study, although CyLens-70B does not match the exact CVE identifier, it effectively aligns with the evidence provided in the threat report and demonstrate a strong threat detection capability.
-This effectiveness is further supported by the results shown in Figure 5-(c)(d).
-RF9: CyLens demonstrates robust threat correlation.
-Table 5 presents the hit ratio for the top 10 correlated threats (CVEs) as reported in MITRE-CVE [14].
-We evaluate three correlation strategies: (i) Direct correlation, (ii) CVEs as intermediates, and (iii) CWEs as intermediates, as described above.
+From the above case study, although CyLens-70B does not match the exact CVE identifier, it effectively aligns with the evidence provided in the threat report and demonstrate a strong threat correlation capability.
+This effectiveness is further supported by the results shown in Figure 5-(c)(d), where CyLens presents stronger correlation relevance w.r.t.
+the correct products and threat categories.
+RF8: CyLens demonstrates robust expansive correlation.
+Table 7 presents the hit ratio for the top 10 expanded threats (CVEs) as reported in MITRE-CVE (Corporation, nd).
+We evaluate three correlation strategies: (i) Direct expansion, (ii) CVEs as intermediates, and (iii) CWEs as intermediates.
 We have corresponding observations:
-Table 5: Evaluation results (HIT@10) on ❹ Correlation task using historical threats. The order of LLM names (left-to-right) follows the same sequence as in Tables 2 (top-down).
+Table 7.
+Evaluation results (HIT@10) on expansive correlation task using historical threats (we do not evaluate on 0-day threats as they have incomplete reports of concurrent CVEs).
+The order of LLM names (left-to-right) follows the same presentation sequence as in Tables 3 (top-down).
 SL LY 4O O1 GM LA DS CD C1 C8 C70
@@
 w/ CVE 7.3 4.5 11.6 25.4 6.7 0.0 4.3 12.5 60.4 79.0 81.2
-w/ CWE 0.0 0.0 0.0 0.8 0.0 0.0 0.0 0.0 41.7 70.8 71.6
+w/ CWE 0.0 0.0 0.0 0.8 0.0 0.0 0.0 0.0 36.2 55.4 60.7
     • (i)
-      While baseline models largely struggle to correlate CVEs, CyLens demonstrates significant effectiveness (e.g., 77.5% Hit@10 with CyLens-70B).
+      While baseline models largely struggle to expansively correlate CVEs, CyLens demonstrates significant effectiveness (e.g., 77.5% HIT@10 with CyLens-70B).
     • (ii)
-      Incorporating CVEs as intermediate evidence improves correlation performance.
+      Incorporating CVEs as intermediate evidence improves expansive correlation.
 However, baseline models tend to propagate erroneous CVEs from the prior detection step, which misguides subsequent threat correlations.
     • (iii)
-      Incorporating CWEs degrades correlation performance due to their inherent ambiguity.
-Since a CWE can broadly relate to many CVEs, it introduces ambiguity, impeding accurate threat correlation.
-Despite this challenge, CyLens remains robust (e.g., 71.6% Hit@10 with CyLens-70B) in incorporating ambiguous intermediate evidence while maintaining strong correlation effectiveness.
-5.4 Threat Prioritization
-Real-world cyber threats often co-occur across cyber events [77, 73], necessitating that CTI teams distinguish threat severity and prioritize subsequent patches and mitigations.
+      Incorporating CWEs degrades expansive correlation due to their inherent ambiguity. Since a CWE can broadly relate to many CVEs, it introduces ambiguity, impeding accurate correlation.
+5.4. Threat Prioritization
+Real-world cyber threats often co-occur across cyber events (Rahman and Williams, 2022; Peng et al., 2020), necessitating that CTI teams to distinguish threat severity and prioritize subsequent patches and mitigations.
 Our evaluation focuses on two prioritization measurements: severity level and analysis of exploitation dynamics.
@@
 Threat severity can be assessed from threat reports even when CVE identifiers are not disclosed or when zero-day threats have not yet been assigned CVEs.
-We evaluate severity using 8 metrics defined in CVSS 3.x [60], which collectively capture a threat’s exploitability and impact.
+We evaluate severity using 8 metrics defined in CVSS 3.x (National Institute of Standards and Technology (2025), NIST), which collectively capture a threat’s exploitability and impact.
 Each severity metric is treated as a classification task.
 For example, the Attack Vector metric has four predefined categories: Network, Adjacent Network, Local, and Physical.
-In cases where a threat report lacks sufficient evidence for CVSS metrics, we introduce an additional category, "N/A", for each metric.
+If a threat report lacks sufficient evidence for CVSS metrics, we introduce an additional label, ”N/A”, for each metric.
 Furthermore, we classify the base severity level from four categories: Low, Medium, High, and Critical.
-This classification aggregates the ratings from 8 CVSS metrics, providing an overall severity assessment aligned with CVSS standards.
-RF10: CyLens maintains stable effectiveness.
+This classification aggregates the ratings from eight CVSS metrics, providing an overall severity assessment aligned with CVSS standards.
+RF9: CyLens maintains stable effectiveness.
 Table 6 shows the accuracy of all models.
 While baselines vary widely across CVSS metrics (e.g., GPT-o1 scores 92.91% on Attack Complexity but only 20.43% on Integrity Impact), CyLens remains consistently strong.
-Notably, CyLens-8B achieves $>$90% accuracy across 8 metrics for historical threats, demonstrating its reliability in handling different severity factors.
-This stability comes from CyLens ’s ability to adapt to different evidence types.
-Each CVSS metric relies on specific context—Privileges Required (PR) depends on access control policies, while Attack Complexity (AC) is influenced by real-world constraints and attacker skill level.
-By effectively integrating diverse evidence sources, CyLens reduces inconsistencies in severity analyses, which aligns with RF4.
-Additionally, we find that zero-day threats are not always harder to analyze, aligning with RF5.
-Some evidence, like some user actions in exploitations, is already known from past threats, facilitating assessment on newly emerging threats.
-We also observe that larger LLMs do not always perform better, consistent with RF6.
-While bigger models tend to provide more detailed reasoning, this can sometimes lead to inaccurate analysis.
-For example, CyLens-70B does not always outperform CyLens-8B across all severity metrics.
+Notably, CyLens-8B achieves $>$85% accuracy across 8 metrics, demonstrating its reliability across different severity factors.
+This stability comes from CyLens ’s adaptability across different evidence types.
+Each CVSS metric relies on specific evidence, for example, Privileges Required (PR) depends on access control policies, while Attack Complexity (AC) is influenced by system constraints and attacker’s skill.
+CyLens reduces inconsistencies in severity analyses by effectively integrating diverse evidence sources.
 Exploitation Dynamics.
 Besides severity analysis using CVSS, we also evaluate CyLens’s effectiveness in analyzing dynamic exploitation behaviors.
-Specifically, we use the EPSS score [29] and conduct two types of analysis: (1) pinpointing historical exploitation at an earlier timestamp and (2) forecasting future exploitation at a later timestamp.
-Figure 6: With $\mathcal{T}_{0}$ as the “current” timestamp (varying among threats), we compute RMSE for historical EPSS scores over the past three months (interpolation) and future EPSS scores for the next three months (prediction) w.r.t.
+Specifically, we use the EPSS score (FIRST.org, 2025) and conduct two types of analysis: (1) pinpointing historical exploitation at an earlier timestamp and (2) forecasting future exploitation at a later timestamp.
+Note that we exclude zero-day threats from the exploitation analysis due to the incompleteness of EPSS data.
+Figure 6.
+With $\mathcal{T}_{0}$ as the “current” timestamp (varying among threats), we compute RMSE for historical EPSS scores over the past three months (interpolation) and future EPSS scores for the next three months (prediction) w.r.t.
 $\mathcal{T}_{0}$.
@@
 To ensure accurate evaluation, we match each EPSS score to the closest available recorded score.
-For example, if an EPSS score is not reported exactly on 10/01/2023, we use the closest available date (e.g., 09/25/2023) from EPSS reports [14] as the ground truth.
-RF11: LLMs vary in interpolation and prediction.
+For example, if an EPSS score is not reported exactly on 10/01/2023, we use the closest available date (e.g., 09/25/2023) from EPSS reports (Corporation, nd) as the ground truth.
+RF10: LLMs vary in interpolation and prediction.
 As shown in Figure 6, LLMs perform slightly better in interpolating historical EPSS scores (lower RMSE) than in predicting future scores.
-However, we also observe that local LLMs, such as SevenLLM, perform obviously worse in both interpolation and prediction due to their static inherent knowledge, which lacks adaptability to evolving cyber threats.
-On the contrary, CyLens, despites its local deployment, performs similar to API-based LLMs (e.g., Gemini), which demonstrates the necessity of equipping LLMs with domain expertise.
+However, we also observe that local LLMs, such as SevenLLM, perform obviously worse in both interpolation and prediction due to their static inherent knowledge, lacking adaptability to the emerging exploitation.
+On the contrary, CyLens performs similar to API-based LLMs (e.g., Gemini), which demonstrates the necessity of equipping LLMs with domain expertise.
 This enables LLMs to adapt to the dynamic nature of CTI tasks.
-RF12: Contextual knowledge is essential. Beyond Figure 6, we observe that some cyber threats undergo dramatic shifts in EPSS scores due to emerging exploitation activities.
-Case Study 5
-For instance, CVE-2024-27956 saw a sharp increase in EPSS on 02/15/2025, jumping from 0.31% to 55.37% [19].
+RF11: Contextual awareness is required.
+We also observe that some cyber threats undergo dramatic shifts in EPSS scores due to emerging exploitation activities.
+This implies the importance of not only fitting to EPSS trends but also incorporating real-world events to enhance CTI capability.
+For instance:
+Case Study 5.
+CVE-2024-27956 presents a sharp increase in EPSS scores on 02/15/2025, jumping from 0.31% to 55.37% (CVE Details, 2024).
 No LLM was able to accurately predict this trend without contextual knowledge of real-world events.
-This implies the importance of not only fitting to EPSS trends but also incorporating real-world evidence in real time to enhance CTI adaptability.
-5.5 Remediation
-Table 6: Evaluation results on ❻ Remediation.
+Figure 7.
+Evaluation results by different groups of threats: (a)–(d) correspond to threats targeting the same product, (e)–(h) to those affecting the same vendor, and (i)–(l) to threats within the same category.
+We further divide these threats into historical and zero-day groups, with dataset statistics provided in Table 4.
+The evaluation metrics used are accuracy for both threat actor attribution and third-party advisory suggestions.
+Abbrevation: ”Bd CTI” – CyLens fine-tuned by broad CTI tasks (same setting as section 5); ”Sp CTI” – CyLens fine-tuned by specialized CTI tasks.
+Additional evaluation results are shown in Figure 8.
+5.5. Remediation
+Table 8.
+Evaluation results on ❻ Remediation.
 We evaluate whether LLMs can correctly generate: (i) patching or mitigation tools – measured by accuracy, (ii) code patches – measured by syntactic similarity, (iii) patching or mitigation methods – measured by BERT Score, and (iv) reference advisories – measured by accuracy.
@@
 Finally, we evaluate CyLens’s capability in generating accurate remediation strategies, focusing on four aspects: suggesting the correct tool for patching or mitigation (e.g., OpenSSL for encryption), generating available code patches (e.g., GitHub commits), proposing appropriate mitigation methodologies (e.g., hardening an OS), and identifying the correct advisory to consult (e.g., NVIDIA Security Advisory).
-RF13: CyLens is effective across all remediation tasks.
-Consistent with RF4 and RF10, we demonstrate CyLens’s effectiveness in handling various remediation objectives as shown in Table 6, including both identification-based (suggesting tools and advisories) and generation-based (producing code patches or mitigation strategies) approaches.
-In contrast, baseline models, lacking CTI-specific expertise, often exhibit imbalanced performance across remediation tasks.
-For example, Gemini performs well in generating code patches but struggles to identify the correct advisories.
-Figure 7: Evaluation results across multiple CTI tasks based on different levels of organizational interests: (a)-(d) product-level, (e)-(h) vendor-level, and (i)-(l) threat category-level.
-The evaluation metrics used are accuracy for threat actor attribution, F1 score for attack infrastructure contextualization, accuracy for CVSS base severity classification, and accuracy for third-party advisory suggestion.
-Model names are SL: SevenLLM, O1: ChatGPT-O1, GM: Gemini-Pro, C1/C8: CyLens-1B/8B.
-RF14: Large-scale expertise is crucial.
-Effective remediation often relies heavily on existing cybersecurity practices [43].
-Therefore, embedding large-scale historical remediation knowledge is crucial when developing LLMs for CTI applications.
-Even for zero-day threats, well-trained LLMs can leverage their inherent knowledge to derive useful insights, improving response effectiveness even in cases where no explicit remediation has been documented.
-5.6 Organizational Customization
-Previous experiments demonstrate CyLens’s capability in analyzing a wide range of cyber threats, providing a strong foundation for addressing specific organizational needs.
-Given that threats of particular interest—such as those associated with specific products, vendors, or threat categories—may lack sufficient volume for fine-tuning [58, 35], it is essential to first develop LLMs with comprehensive CTI analysis capabilities.
-These LLMs can then be incrementally customized to align with specific organizational needs.
-We categorize organizational needs into three levels:
-    • Product-Level – Focusing on cyber threats affecting specific products, a common approach among security teams in large enterprises managing product-specific CTI [21].
-    • Vendor-Level – Targeting threats that impact an entire vendor (i.e., company), a perspective often adopted by small and medium-sized businesses or industry groups [67].
-    • Category-Level –Concentrating on specific threat categories, which is particularly relevant to research institutions conducting specialized cybersecurity studies [49].
-Settings.
-To cover diverse organizational needs, we collect threats associated with (1) specific products (Acrobat, Chrome, or Fedora), (2) specific vendors (openSUSE, GitLab, or Dell), and (3) specific threat categories (Overflow, DoS, or XSS).
-Thus, we collect a total of nine groups of threats, with each group comprising more than 1,000 threat reports.
-We allocate 50% of these threat reports—sourced from third-party references in the MITRE-CVE [15], NVD, and Exploit-DB [62] websites [59]—for independently fine-tuning CyLens-1B/8B, while reserving the remaining for evaluation.
-We compare CyLens to the SOTA model, SevenLLM, and two industry-leading LLMs, GPT-o1 and Gemini.
-Figure 7 presents the results across multiple CTI tasks.
-RF15: CyLens is adaptable to diverse organizational needs.
-From Figure 7, we observe that CyLens maintains high effectiveness even with independent customization (fine-tuning) across different groups of cyber threats.
-For instance, CyLens-8B achieves nearly 100% accuracy in providing remediation advisories for various vendors and threat categories.
-This result further demonstrates its adaptability despite the diverse nature of threats across products, vendors, and categories—factors that significantly impact baseline LLMs, causing their performance to fluctuate.
-The finding implies CyLens ’s superior capability to generalize while maintaining high effectiveness in tailored needs.
-6 Discussion
-6.1 Ablation Study
+RF12: Broad expertise is crucial.
+Consistent with RF1 and RF9, we demonstrate CyLens’s effectiveness in addressing a wide range of remediation objectives, as shown in Table 8.
+These include both identification-based tasks (e.g., suggesting appropriate tools and advisories) and generation-based tasks (e.g., producing code patches or mitigation strategies).
+In contrast, baseline models often exhibit imbalanced performance across remediation tasks.
+For instance, while Gemini-Pro performs well in generating code patches, it struggles to correctly identify relevant advisories.
+We attribute this disparity to the fact that effective remediation often depends on familiarity with established cybersecurity practices (Jacobs et al., 2020).
+As such, embedding large-scale historical remediation knowledge is essential when developing LLMs for CTI applications.
+Notably, even for zero-day threats, a well-trained model like CyLens can leverage its inherent domain knowledge to derive meaningful insights, thereby enhancing response effectiveness in scenarios where no explicit remediation guidance exists.
+6. Evaluation Results: Specialized CTI
+Despite CyLens’s scalability in capturing a broad threat landscape, real-world CTI often involves targeted interests (Mvula et al., 2024; Optiv, 2025).
+In this section, we demonstrate CyLens’s adaptability in addressing three specialized groups of threats:
+    • Product – Focusing on cyber threats affecting specific products, a common scenario among security teams in large enterprises managing product-specific CTI (Cybersecurity and Infrastructure Security Agency, 2023).
+    • Vendor – Targeting threats that impact an entire vendor (i.e., company), which is a common practice among industry groups, particularly those have focused products or services.(Optiv, 2025).
+    • Category – Concentrating on specific threat categories, which is particularly relevant to research institutions conducting specialized cybersecurity studies (Kim and Kwon, 2022).
+Specialized Instruction Tuning.
+Following the same re-training process described in section 3.3, CyLens is independently fine-tuned on threats associated with specific products (MacOS, Chrome, or Linux Kernel), vendors (IBM, Adobe, or Dell), and categories (Overflow, DoS, or XSS).
+These selections are based on their exploitation popularity as recorded in the MITRE-CVE database (Corporation, 1999).
+The fine-tuning datasets are derived from a subset of the original collections (Table 11) that include only historical threats, with no overlap with the evaluation dataset (Table 4).
+Evaluation Targets.
+As detailed in section 4, we evaluate CyLens using real-world threat reports collected from multiple platforms (Oracle Corporation, 2025; Red Hat, Inc., 2025; Adobe Systems Incorporated, 2025; Apache Software Foundation, 2025a; Debian Project, 2025; openSUSE Project, 2025; Fedora Project, 2025).
+From these sources, we extract threats targeting the selected products, vendors, and categories.
+The corresponding dataset statistics are reported in Table 4.
+We compare the specialized CyLens-8B model with two SOTA, SevenLLM and Lily-Cyber, and two industry-leading LLMs, GPT-o1 and Gemini.
+Additionally, we include a non-specialized CyLens-8B fine-tuned on a broad range of CTI tasks under the same settings as in section 5.
+The evaluation results across multiple CTI tasks are summarized in Figure 7 and 8.
+We have the following findings:
+RF13: CyLens is adaptable to diverse specialized needs.
+From Figure 7, we observe that CyLens maintains high effectiveness even with independent instruction-tuning across different groups of cyber threats.
+For instance, CyLens-8B achieves nearly 100% accuracy in providing remediation advisories for Adobe.
+This result further demonstrates CyLens’s adaptability despite the diverse nature of threats across products, vendors, and categories—factors that significantly impact baseline LLMs, causing their performance to fluctuate.
+RF14: Data scarcity matters.
+Cyber threats targeting specific products or even entire vendors often suffer from data scarcity, particularly a lack of historical threats that are important for experience-based remediation.
+This data scarcity can hinder the development of countermeasures, especially for zero-day threats.
+For example:
+Case Study 6.
+As of 2024, only 9 cyber threats related to Open Redirection vulnerabilities had been reported for Dell.
+However, within the first three months of 2025, 2 new threats emerged—targeting Dell Unity (CVE-2025-24381) and Networker (CVE-2025-21104)—highlighting the challenge of generating timely and effective mitigation strategies with limited historical data.
+RF15: A broad view of the threat landscape is necessary.
+When developing CTI solutions for specific products or vendors, relying solely on threat data limited to those entities is often insufficient (as highlighted in RF14).
+It is essential to consider the broader threat landscape to uncover relevant analytical patterns or remediation strategies.
+This is demonstrated by the effectiveness of CyLens —both CyLens-1B and CyLens-8B achieve significantly higher accuracy in advisory suggestion for Dell-related threats (¿70%) compared to all baseline methods (¡50%).
+A representative case study is:
+Case Study 7.
+In studying the remediation of CVE-2025-21104, CyLens-8B successfully identifies that reconfiguring the web server to enforce strict parsing rules and ignore unexpected Host header values can mitigate the vulnerability—an approach typically incorporated in newer system versions.
+This insight aligns with Dell’s official advisory DSA-2025-124, which includes a security update for the Dell NetWorker Management Console addressing the issue.
+7. Discussion
+7.1. Ablation Study
 We conduct ablation studies to assess the impact of different components within CyLens.
-Table 7 presents the results, where we independently exclude pre-training (PT), fine-tuning (FT), and four optional modules: topic modeling (TOM), named entity recognition (NER), relation extraction (REL), and retrieval-augmented generation (RAG).
-Notably, reasoning (REA) and summarization (SUM) cannot be excluded, as they are essential steps in text generation.
-Table 7: Ablation study on CyLens-8B by selecting four representative CTI tasks: (1) attributing threat actors (measured by accuracy), (2) contextualizing attack infrastructure (measured by F1 score), (3) classifying CVSS base severity levels (measured by accuracy), and (4) suggesting advisories for remediation (measured by accuracy).
+Table 9 presents the results, where we independently exclude re-training (RT), fine-tuning (FT), and four functional modules: TOM, NER, REL, and RAG.
+Notably, REA and SUM cannot be excluded, as they are essential inference steps.
+Table 9.
+Ablation study on CyLens-8B by selecting four representative CTI tasks: (1) attributing threat actors (measured by accuracy), (2) contextualizing attack infrastructure (measured by F1 score), (3) classifying CVSS base severity levels (measured by accuracy), and (4) suggesting advisories for remediation (measured by accuracy).
 Ablation Threat Actor Attack Infra CVSS Base Advisory
 Hist. 0-Day Hist. 0-Day Hist. 0-Day Hist. 0-Day
-w/o PT 66.22 62.31 70.05 61.44 53.62 45.90 57.25 54.36
+w/o RT 66.22 62.31 70.05 61.44 53.62 45.90 57.25 54.36
 w/o FT 82.42 70.15 86.47 78.31 86.48 61.75 82.56 77.23
@@
 w/o RAG 84.17 80.26 86.53 82.71 95.23 89.64 85.14 83.42
-We observe that pre-training has the most significant impact, as it embeds massive CTI knowledgeinto the model.
+We observe that re-training has the most significant impact, as it embeds massive CTI knowledgeinto the model.
 While fine-tuning is more lightweight, its influence remains non-negligible in refining CyLens ’s CTI capabilities.
@@
 These findings reinforce the necessity of incorporating all proposed techniques to achieve an effective and well-rounded system design for CTI applications.
-6.2 Adapting to Computational Constraints
-Similar to Meta’s Llama series, which offers models at different scales [56, 57, 92], CyLens is available in three sizes: 1B, 8B, and 70B, allowing it to accommodate varying computational constraints.
-While CyLens-1B is less effective compared to the 8B and 70B versions, its lightweight design is crucial for resource-limited deployments.
+7.2. Adapting to Computational Constraints
+Similar to Meta’s Llama series, which offers models at different scales (Meta-llama, 2024a, b; Touvron et al., 2023), CyLens is available in three sizes: 1B, 8B, and 70B, allowing it to accommodate varying computational constraints.
+While CyLens-1B is less effective compared to the 8B and 70B, its lightweight design is crucial for resource-limited deployments.
 In scenarios where computational resources are highly constrained, having a functional LLM is often more practical than striving for state-of-the-art performance.
 Notably, despite its smaller size, CyLens-1B still outperforms industry-leading LLMs in most CTI tasks, reinforcing its value as an efficient yet capable model for cybersecurity applications.
-6.3 Privacy-Preserving Deployment
-A key advantage of customizing local LLMs is enhanced privacy preservation.
-Unlike API-based services such as OpenAI’s GPT series, which inherently risk exposing user intentions and confidential data, local models ensure that sensitive information remains under organizational control.
-API-based solutions may not only transmit proprietary data to external servers but could also use it for model training and further development, raising concerns about data security and regulatory compliance [65, 66, 80, 99, 113].
-By developing localized, task-specific LLMs, organizations gain full control over model behavior, customization, and data security.
-Such privacy-preserving deployment is particularly critical in high-stakes fields like CTI, where threat intelligence, incident reports, and security strategies must remain confidential to prevent adversarial exploitation and ensure compliance with data protection regulations.
-6.4 Beyond Specific Needs
-Real-world cyber threats are rarely confined to specific products, vendors, or categories [77].
-Instead, they manifest within a complex and interconnected ecosystem where vulnerabilities, exploits, and attack methodologies transcend traditional boundaries.
-This interdependence underscores the necessity of developing CTI solutions that go beyond addressing specific organizational needs (e.g., tailored threat intelligence) and are also capable of capturing the broader, multifaceted threat landscape.
-7 Conclusion
-This paper introduced CyLens, a LLM-powered copilot for cyber threat intelligence.
-CyLens streamlines the entire threat management lifecycle, encompassing tasks from threat attribution to remediation.
-It assimilates knowledge from a vast number of threat reports and utilizes a retrieval module for real-time information, enabling it to adeptly handle both historical and zero-day threats.
-Rigorous evaluations have demonstrated that CyLens consistently outperforms industry-leading LLMs and SOTA agents across multiple CTI tasks.
-Its adaptability shines through its customization capabilities, which can be tailored to specific threat categories or organizational requirements.
-Additionally, the release of its training datasets fosters further research in the field.
+8. Conclusion
+We introduce CyLens, a LLM-powered agentic CTI system.
+CyLens supports the entire threat management lifecycle, encompassing tasks from threat attribution to remediation.
+It accumulates CTI knowledge from massive threat reports and utilizes a functional module for real-time analysis, enabling it to adeptly handle both historical and zero-day threats.
+Intensive evaluations have demonstrated the scalability and adaptability of CyLens across massive CTI tasks and tailored needs of threat management.
+By consistently outperforming industry-leading LLMs and SOTA agents, CyLens offers scalable and adaptive solution and serves as a blueprint for integrating agentic LLMs into CTI practice.
 Appendix A Pre-Training: Complementary Information
-Table 8: The data sources we utilize and the corresponding evidence metadata collected from each database.
+Table 10. The data sources we utilize and the corresponding evidence metadata collected from each database.
 CTI Task Analysis Target Source Database
@@
 TTPs
-Exploit Path
+Campaign
 Affected System
@@
 CWE Identifier
-CVE Identifier
 CVSS Metrics
@@
 Advisory
-Table 9: Base LLM assignment and settings of CyLens’s development, including pre-training and fine-tuning. “Cp” – Corpus; “Tk” – token.
-CyLens Base LLM Pre-Training
-# Cp # Tk $T$ $T_{1}$ $T_{2}$ $\beta$
-CyLens-1B Llama-3.2-1B 28K 0.27B 4 3 3 1.0
-CyLens-8B Llama-3.1-8B 115K 1.1B 10 5 5 1.0
-CyLens-70B Llama-3.3-70B 271K 2.5B 20 10 10 1.0
-A.1 Data Source and Metadata Collection
-Complementary to section 3.1, Table 8 provides a detailed overview of the cyber threat databases we utilize, along with the corresponding metadata collected for various CTI tasks (❶-❻) and analytical targets.
+A.1. Data Source and Metadata Collection
+Complementary to section 3.2, Table 10 provides a detailed overview of the cyber threat databases we utilize, along with the corresponding metadata collected for various CTI tasks (❶-❺) and analytical targets.
 MITRE-CVE and NVD.
-Our collected data is primarily centered around CVEs, leveraging the MITRE-CVE [14] and NVD [59] databases, which serve as well-structured repositories of publicly disclosed vulnerabilities.
+Our collected data is primarily centered around CVEs, leveraging the MITRE-CVE (Corporation, nd) and NVD (National Institute of Standards and Technology, 2025) databases, which serve as well-structured repositories of publicly disclosed vulnerabilities.
 These databases provide foundational information, including CVE identifiers, descriptions, severity scores, affected products, and references to external sources.
 Exploit-DB.
-To further enrich our dataset and enhance contextual understanding, we incorporate additional cyber threat intelligence (CTI) sources, including Exploit-DB [62].
+To further enrich our dataset and enhance contextual understanding, we incorporate additional cyber threat intelligence (CTI) sources, including Exploit-DB (Offensive Security, 2025).
 Exploit-DB provides reports on security advisories and threat intelligence insights, complementing the structured data from MITRE-CVE and NVD.
 Notably, Exploit-DB is also CVE-centric; however, unlike MITRE-CVE and NVD, it lacks structured connections to additional evidence.
-Therefore, our data collection primarily focuses on extracting raw threat descriptions and attack vector classifications (e.g., "Local") to supplement our dataset with additional contextual details.
+Therefore, our data collection primarily focuses on extracting raw threat descriptions and attack vector classifications (e.g., ”Local”) to supplement our dataset with additional contextual details.
 Other MITRE Databases.
 Beyond these primary databases, we integrate metadata from additional sources to enrich our dataset with diverse perspectives on cyber threats.
-Specifically, we leverage CWE [16] to systematically categorize software weaknesses, providing a structured foundation for understanding vulnerability root causes.
-CAPEC [17] offers a comprehensive taxonomy of attack patterns, enabling us to analyze how adversaries exploit known weaknesses in real-world scenarios.
-MITRE ATT&CK [18] serves as a crucial resource for mapping adversary tactics, techniques, and procedures (TTPs), facilitating in-depth attribution and behavioral analysis.
+Specifically, we leverage CWE (Corporation, 2006) to systematically categorize software weaknesses, providing a structured foundation for understanding vulnerability root causes.
+CAPEC (Corporation, 2007) offers a comprehensive taxonomy of attack patterns, enabling us to analyze how adversaries exploit known weaknesses in real-world scenarios.
+MITRE ATT&CK (Corporation, 2013) serves as a crucial resource for mapping adversary tactics, techniques, and procedures (TTPs), facilitating in-depth attribution and behavioral analysis.
 Third-Party References.
@@
 Among them, GitHub commits to serve as a valuable source for tracking vulnerability disclosures, exploit development, and security patch discussions.
-Apache security reports [5] provide detailed documentation on vulnerabilities and remediation efforts specific to widely used web and enterprise software.
-Meanwhile, IBM X-Force threat intelligence [39] contributes curated threat reports, real-world attack observations, and insights into evolving threat landscapes.
+Apache security reports (Apache Software Foundation, 2025b) provide detailed documentation on vulnerabilities and remediation efforts specific to widely used web and enterprise software.
+Meanwhile, IBM X-Force threat intelligence (IBM X-Force, 2025) contributes curated threat reports, real-world attack observations, and insights into evolving threat landscapes.
 The collected metadata, detailed in Table 10, enhances the comprehensiveness of our dataset by offering multifaceted insights into vulnerability exploitation patterns, adversarial attack techniques, and effective mitigation strategies.
 By integrating these diverse sources, we enable a richer and more contextualized analysis of cyber threats, supporting improved detection, attribution, and defense mechanisms.
-A.2 Threat Corpus Generation Details
+A.2. Threat Corpus Generation Details
 Complementary to section 3.2, we detail the prompts used for generating and refining the threat corpus, specifically in Prompt 1 and 2.
 During output sampling, we adjust the temperature and top-p parameters within a range of 0.2 to 1.0 to balance creativity and determinism.
-An example of a generated threat report is provided in Example 4.
+An example of a generated threat report is provided in Example A.1.
 Additionally, we present the statistical overview of the corpus used for pre-training in Table 1.
-Prompt 1
+Prompt 1.
 Below is the prompt template used for threat corpus generation.
 […]
-Prompt 2
+Prompt 2.
 Below is the prompt used for threat corpus revision.
@@
 $\cdots$ Report to Revise: [Insert Cyber Threat Report Here] Output:
-Example 4
+Example 0.
 Below is an example of threat report.
@@
 The lack of sandboxing further exacerbates the issue by allowing the application broader access to system resources than necessary, potentially exposing additional data $\cdots$ Mitigation: Users are advised to update the $\cdots$
-Appendix B Fine-Tuning: Complementary Information
-Complementary to section 4.1, we present the statistics of the generated fine-tuning dataset in Table 10.
+Appendix B Instruction Tuning: Complementary Detail
+Complementary to section 3.4, we present the statistics of the generated instruction-tuning dataset in Table 11.
 Notably, when generating each data instance using a cascading reasoning layout, not all analytical targets are included due to variations in their availability.
-Table 10: Statistics of the fine-tuning dataset with a cascading reasoning layout, where each analytical target appears in varying volumes of data.
-Additionally, different scales of fine-tuning data are used for different CyLens models to align with their respective capacities and optimization needs.
+Table 11.
+Statistics of the instruction-tuning dataset with a cascading reasoning layout, where each analytical target appears in varying volumes of data.
+Additionally, different scales of data are used for different CyLens models to align with their respective capacities and optimization needs.
 CTI Task Analytical Target # Sample # Fine-Tuning Samples
 1B 8B 70B
-Threat Actor 37.6K 0.8K 11.5K 37.6K
-TTPs 35.5K
-Exploit Path 21.8K
-Affected System 37.6K
-Attack Infra 28.3K
-Attack Impact 17.6K
-CVE Identifier 37.6K
-CWE Identifier 37.6K
-CVE Identifier 22.9K
-CVSS Metrics 37.6K
-EPSS Records 33.1K
-Tool Use 15.3K
-Code Patch 8.2K
-Methodology 32.4K
-Advisory 27.5K
+Threat Actor 37.6K 1.2K 11.5K 37.6K
+TTPs 35.5K 1.2K 11.3K 35.5 K
+Campaign 21.8K 0.8K 7.9K 21.8K
+Affected System 37.6K 1.2K 11.5K 37.6K
+Attack Infra 28.3K 0.9K 10.2K 28.3K
+Attack Impact 17.6K 0.6K 5.4K 17.6K
+CVE Identifier 37.6K 1.2K 11.5K 37.6K
+CWE Identifier 37.6K 1.2K 11.5K 37.6K
+CVSS Metrics 37.6K 1.2K 11.5K 37.6K
+EPSS Records 33.1K 1.3K 12.6K 33.1K
+Tool Use 15.3K 0.7K 4.8K 15.3K
+Code Patch 8.2K 1.1K 3.1K 8.2K
+Methodology 32.4K 1.0K 14.8K 32.4K
+Advisory 27.5K 0.8K 9.6K 27.5K
 Furthermore, CyLens-1B, CyLens-8B, and CyLens-70B utilize different volumes of fine-tuning data, reflecting their respective model capacities and optimization requirements.
-For fine-tuning CyLens-1B, we select 0.8K samples that contain all analytical targets.
+For fine-tuning CyLens-1B, we select a smaller volume of samples that contain all analytical targets.
 In contrast, for CyLens-8B and CyLens-70B, we use a progressively larger dataset, with CyLens-8B utilizing a partial subset and CyLens-70B leveraging the full dataset, despite the potential absence of certain analytical targets in some instances.
-Appendix C Evaluation: Complementary Information
-C.1 Evaluation Data
+Appendix C Inference Module: Complementary Detail
+This section detail the implementation of six NLP modules as described in section 3.5.
+C.1. Topic Modeling
+To perform topic modeling, we leverage BERTopic, a SOTA topic modeling framework that integrates transformer-based embeddings with clustering and class-based TF-IDF.
+In our implementation, we design a two-stage prompt-based workflow to integrate BERTopic with CyLens for dynamic topic modeling.
+In the first stage, CyLens is prompted to assess whether topic modeling is needed based on the input text’s characteristics.
+CyLens receives instructions:
+Prompt 3.
+Necessity of Topic Modeling.
+Given the input query, determine whether topic modeling is appropriate. Hint: topic modeling is suitable when the texts cover multiple themes or the user seeks to explore underlying topics.
+CyLens evaluates the semantic content and returns a binary decision along with reasoning.
+Once CyLens determines that topic modeling is appropriate, it proceeds to construct a structured call to BERTopic via a second-stage prompt.
+This prompt guides CyLens to assemble the necessary inputs for BERTopic and invoke it accordingly.
+The prompt is:
+Prompt 4.
+Calling Topic Modeling.
+You have identified that topic modeling is required.
+Use BERTopic to extract topics from the following documents.
+For each topic, return a descriptive label, a list of representative keywords, and top exemplar documents.
+CyLens structures this as a function call or prompt completion depending on the system design.
+BERTopic is then executed with the provided documents, and the results are passed back to CyLens, which can further head to the next NLP module.
+C.2. Named Entity Recognition (NER)
+To perform NER, we adopt a prompt-based inference approach in CyLens in place of traditional sequence labeling methods.
+Rather than fine-tuning, CyLens is queried at runtime with carefully crafted prompts to identify and categorize named entities directly from the input text.
+The prompt is
+Prompt 5.
+NER Prompt (An Example for Attribution Task).
+System Prompt: You are a cybersecurity threat intelligence assistant specialized in named entity recognition (NER).
+Your task is to extract and categorize entities relevant to threat attribution from unstructured text.
+Focus on identifying entities that help answer the question ”Who is responsible for the attack?” and ”How was the attack carried out?” Instructions: Given a cybersecurity-related document or report excerpt, extract all relevant named entities and classify them into the following categories: Threat Actor: Names or aliases of individuals or groups suspected or known to be behind the activity.
+$\cdots$ Return the results in the following structured JSON format.
+CyLens then returns a list of entities, each annotated with its type and surface form.
+C.3. Relation Extraction
+To extract semantic relationships between entities in cybersecurity texts, we implement relation extraction through prompt-based inference in CyLens.
+This method allows us to bypass traditional supervised training by leveraging the CyLens ’s few-shot and zero-shot reasoning capabilities.
+At inference time, we provide the CyLens with structured prompts that instruct it to identify and describe relationships between extracted entities, such as connections between threat actors and malware, or between attack campaigns and their target sectors.
+A typical prompt format is
+Prompt 6.
+Relation Extraction Prompt
+System Prompt: You are a cybersecurity threat intelligence reasoning assistant.
+Your task is to perform relation extraction to identify and structure meaningful semantic relationships between entities mentioned in cybersecurity texts, particularly in the context of threat attribution.
+The goal is to infer who did what to whom, how, when, and using what tools or infrastructure.
+Instructions: Given a cybersecurity-related passage (such as a threat report or incident summary) and its extracted named entities, identify relevant relationships between entities and represent them as a list of (subject, relation, object) triples.
+Focus on attribution-relevant relations.
+Use only clearly supported relationships based on the input.
+If no relationships are found, return an empty list.
+Return the results in this structured JSON format.
+C.4. Retrieval-Augmented Generation (RAG)
+To enhance the quality of CyLens inference, we implement a RAG framework that integrates external web-based APIs during inference.
+This bridges generative reasoning with real-time access to structured and unstructured data sources, such as open-source intelligence feeds, public CTI repositories, and search APIs.
+The RAG workflow begins with the CyLens analyzing the user’s input to determine whether external retrieval is necessary.
+For example, if the input question involves current threat actor activity, recent campaign data, or infrastructure details (e.g., domains, malware hashes), the system triggers a retrieval phase.
+In this phase, CyLens is prompted to generate a structured search query that captures the essential information needs of the input.
+The prompt format follows a declarative schema, such as: “Formulate a concise search query to retrieve current information about [topic], including actors, tools, targets, and known affiliations.” The generated query is sent to a relevant external data source via API—for instance, querying public CTI portals (e.g., MITRE ATT&CK, VirusTotal, Shodan, or threat feed aggregators) or web search APIs (e.g., Bing Search or Google Custom Search).
+Upon receiving the retrieved content (typically one or more passages or structured documents), the system constructs a new input prompt to CyLens, embedding the original question and the retrieved evidence into the context window.
+This second-stage prompt is crafted using templates such as: “Given the following query and retrieved documents, generate a comprehensive, evidence-based answer.
+Use only the provided content.” This constraint guides CyLens to ground its reasoning in the externally retrieved information, thereby improving factual accuracy and traceability.
+C.5. Reasoning
+To support advanced analytical capabilities in CyLens, we implement modular reasoning and summarization components using prompt engineering.
+Rather than relying solely on fine-tuned models for specific downstream tasks, our architecture employs LLMs (i.e., CyLens itself) as general-purpose reasoning engines, guided by carefully constructed prompts.
+This allows CyLens to remain flexible, easily extensible, and capable of performing domain-specific tasks such as cyber threat correlation, threat actor profiling, and intelligence synthesis without retraining.
+The reasoning module is designed to interpret and connect extracted entities, attributes, and contextual cues across threat intelligence documents.
+We use prompt engineering to simulate multi-step analytical thinking, instructing the LLM to infer relationships, fill in missing links, and assess causality.
+For instance, prompts are constructed as: “Given the following entities and attack context, identify the most plausible attribution and justify your reasoning step-by-step.” This encourages the model to break down the logic behind associations, such as linking a threat actor to a campaign based on observed TTPs, tool reuse, or geopolitical motivation.
+We also use chain-of-thought prompting and structured templates (e.g., tabular reasoning) to improve traceability and consistency of the model’s analytical conclusions.
+C.6. Summarization
+The summarization module in CyLens is responsible for generating high-level, human-readable briefings from dense technical threat reports or multi-source intelligence inputs.
+To achieve this, we implement domain-adapted summarization prompts that guide CyLens to preserve critical cybersecurity-relevant information such as named entities, TTPs, indicators of compromise (IOCs), and timeline markers.
+The prompt format typically follows: “Summarize the following document into a concise threat intelligence report, highlighting threat actors, attack vectors, affected sectors, and key dates.” Depending on the use case, we support multiple summarization styles, including executive summaries, technical briefs, and IOC-centric extracts.
+We also integrate few-shot examples when needed to adapt the tone or structure to a specific user role (e.g., SOC analyst vs.
+CISO).
+Appendix D Evaluation: Complementary Information
+Figure 8. Additional evaluation results across on attack infrastructure contextualization (F1 score) and CVSS base severity classification (accuracy), complemented to Figure 7.
+D.1. Evaluation Data
 Rationale.
@@
 This dual approach is essential for developing resilient cybersecurity frameworks that balance historical insights with forward-looking threat intelligence.
-Table 11: Statistics of evaluation dataset
-Historical Zero-Day
-# Threats 10,000 5,526
-Table 11 presents the volume of our evaluation datasets. We ensure that no threats included in the evaluation appear in the training data.
-C.2 Evaluation Metrics
+D.2. Evaluation Metrics
 We define the evaluation metrics used for each analytical target, along with the strategies and rationale behind their selection.
@@
           Intersection over Union (IoU) – Quantifies the overlap between the predicted and reference TTP sets.
-    • Exploitation Path: Since exploitation paths are usually descriptive observations, we use BERT Score to measure text similarity between the generated and reference descriptions.
+    • Campaign: Since campaign is usually descriptive report, we use accuracy to have a binary judgment of whether LLM-generated campaign meets the reference.
+We also use BERT Score to measure text similarity between the generated and reference descriptions.
     • Affected System & Attack Infrastructure: Similar to TTPs, both affected systems and attack infrastructures are structured as itemized lists.
@@
     • Code Patch & Mitigation: Since both tasks involve generating textual recommendations, we use BERT Score to assess the similarity between the LLM-generated output and the reference.
+D.3. Additional Evaluation Results
+Complementing section 6, Figure 8 presents additional evaluation results on specialized CTI tasks, focusing on attack infrastructure contextualization (measured by F1 score) and CVSS base severity classification (measured by accuracy).
+The results are reported across different products, vendors, and threat categories.